### v1

# DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model

Abstract.
All-in-One image restoration aims to address multiple image degradation problems using a single model, significantly reducing training costs and deployment complexity compared to traditional methods that design dedicated models for each degradation type. Existing approaches typically rely on Degradation-specific models or coarse-grained degradation prompts to guide image restoration. However, they lack fine-grained modeling of degradation information and face limitations in balancing multi-task conflicts. To overcome these limitations, we propose DPMambaIR, a novel All-in-One image restoration framework. By integrating a Degradation-Aware Prompt State Space Model (DP-SSM) and a High-Frequency Enhancement Block (HEB), DPMambaIR enables fine-grained modeling of complex degradation information and efficient global integration, while mitigating the loss of high-frequency details caused by task competition. Specifically, the DP-SSM utilizes a pre-trained degradation extractor to capture fine-grained degradation features and dynamically incorporates them into the state space modeling process, enhancing the model’s adaptability to diverse degradation types. Concurrently, the HEB supplements high-frequency information, effectively addressing the loss of critical details, such as edges and textures, in multi-task image restoration scenarios. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, with 27.69dB and 0.893 in PSNR and SSIM, respectively. These results highlight the potential and superiority of DPMambaIR as a unified solution for All-in-One image restoration.
All-in-One Image Restoration, State Space Model, Prompt Learning
1. Introduction
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Recent advances in deep learning have significantly improved performance in Degradation-specific restoration problems, such as image denoising (Kim et al., 2024; Zeng et al., 2024; Ryou et al., 2024), deblurring (Chen et al., 2024b; Wu et al., 2024a; Mao et al., 2024), deraining(Chen et al., 2024a, 2023; Ba et al., 2022; Peng et al., 2020, 2021; He et al., 2024b), dehazing(Zhang et al., 2024e; Wang et al., 2024c; Zheng et al., 2023; Wu et al., 2023) and low-light enhancement(Ma et al., 2022; Cai et al., 2023; Hai et al., 2023; Wang et al., 2022, 2023b). However, traditional methods typically rely on Degradation-specific models explicitly tailored to individual degradation types. These approaches often require prior knowledge of the degradation type, which limits their practicality in real-world scenarios, such as autonomous driving and nighttime surveillance, where diverse and unknown degradations frequently coexist.
Figure 1. Comparison with SOTA All-in-One Image Restoration Methods on mixed Dataset. Our DPMambaIR achieves consistently better performance.
To address this limitation, recent studies explore All-in-One image restoration frameworks that unify multiple restoration tasks into a single model. Existing methods can be broadly categorized into Mixture-of-Experts (MoE)-based and Prompt-based approaches. MoE-based methods, such as MoFME (Zhang et al., 2024b), dynamically route tasks to specialized experts using an uncertainty-aware router, enabling effective learning of diverse expert functions. However, these methods suffer from high computational costs, increased inference complexity as the number of experts grows, and lack explicit degradation modeling, limiting adaptability to unseen or complex degradations. Prompt-based methods, like OneRestore (Guo et al., 2024a), guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms for flexible and adaptive strategies. While effective, these methods rely on coarse prompts that fail to capture nuanced degradation characteristics and treat prompts as static inputs, limiting their ability to adapt to varying degradation intensities or specific requirements, leading to suboptimal results.
While these approaches reduce computational cost and deployment complexity by leveraging feature sharing and multi-task learning, two key challenges remain unresolved. First, existing models lack fine-grained degradation awareness and often rely on coarse degradation labels to guide the restoration process. For example, images with varying degrees of haze are typically assigned a single haze label, failing to capture the nuanced variations in degradation levels. This simplification leads to suboptimal restoration performance, as the model cannot adapt to the diverse characteristics of different degradations. Second, current methods struggle to balance the competing requirements of varying restoration tasks. For instance, training low-light enhancement, which focuses on restoring low-frequency contrast information, alongside deblurring, which prioritizes the recovery of fine local details, often results in task competition. This inherent task competition makes it challenging to optimize a single network for all restoration tasks simultaneously and remains a fundamental challenge for All-in-One image restoration frameworks.
To address these challenges, we propose a novel All-in-One image restoration framework, DPMambaIR. The framework is built upon a Degradation-aware Prompt State Space Model (DP-SSM) and a lightweight High-frequency Enhancement Block (HEB). Specifically, DP-SSM employs a pre-trained degradation extractor to extract degradation embeddings, which are used to modulate the state transition, input, and output matrices of the state space equations. This enables fine-grained degradation modeling and efficient global degradation integration. Meanwhile, HEB mitigates the loss of high-frequency details caused by task competition among different restoration tasks. Extensive experiments on a mixed dataset containing seven degradation types demonstrate the superiority of DPMambaIR. As shown in Figure 1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore. Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, showcasing its potential as an effective and robust All-in-One image restoration solution.
The contributions of this paper can be summarized:
  • $\bullet$
    We propose a novel degradation extractor capable of capturing fine-grained degradation features from complex degraded images. These features provide critical priors to improve the performance of All-in-One image restoration.
  • $\bullet$
    We propose a Degradation-aware Prompt State Space Model (DP-SSM) that incorporates fine-grained degradation features into the state space modeling process, enabling both global degradation modeling and fine-grained degradation awareness to effectively handle diverse degradation types within a unified framework.
  • $\bullet$
    Extensive experiments on a mixed dataset containing seven types of degradation demonstrate that DPMambaIR achieves state-of-the-art performance on PSNR and SSIM metrics, outperforming existing methods. The results validate the effectiveness and superiority of the proposed approach.
Figure 2. The overall framework of DPMambaIR consists of a shallow feature extractor and multiple DPSS blocks, with an upsampler implemented by pixel shuffle and a downsampler by pixel unshuffle. Additionally, a degradation extractor is included to extract degradation embeddings, which are then inserted into the DPSS blocks.
2. Related Work
2.1. All-in-One Image Restoration
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Traditional methods, such as Dark Channel Prior (He et al., 2010) and Color Line Prior (Fattal, 2014), rely on handcrafted priors to model degradation but struggle with complex scenarios due to predefined assumptions. The rise of deep learning (LeCun et al., 2015) has significantly advanced tasks like such as image super-resolution(Peng et al., 2024c, d, 2025c; Wang et al., 2024b; Saharia et al., 2022; Gao et al., 2023), image denoising (Kim et al., 2024; Zeng et al., 2024; Ryou et al., 2024), deblurring (Chen et al., 2024b; Wu et al., 2024a; Mao et al., 2024), deraining(Chen et al., 2024a; Pan et al., 2023; Chen et al., 2023; Ba et al., 2022; Peng et al., 2020, 2021; He et al., 2024b), dehazing(Zhang et al., 2024e; Wang et al., 2024c; Zheng et al., 2023; Wu et al., 2023), desnowing(Guo et al., 2025; Chen et al., 2021; Zhang et al., 2021) and low-light enhancement(Ma et al., 2022; Cai et al., 2023; Hai et al., 2023; Wang et al., 2022, 2023b; Ma et al., 2024). These approaches, leveraging convolutional neural networks (CNN)-based (LeCun et al., 1989), transformer-based (Vaswani et al., 2017; Pan et al., 2025a, 2024), or diffusion-based (Ho et al., 2020) architectures, enable powerful feature learning for image restoration. However, most existing works focus on addressing a single type of degradation, which limits their applicability in real-world scenarios where images often suffer from diverse and unknown types of degradation.
Recent advancements in All-in-One image restoration frameworks focus on enhancing flexibility and efficiency by leveraging various multi-task learning techniques such as Mixture-of-Experts (MoE) and text-based or vision-based prompts to handle diverse restoration tasks. For example, AirNet(Li et al., 2022) employs contrastive learning to extract degradation representations for guiding restoration. PromptIR(Potlapalli et al., 2023) integrates end-to-end prompt generation and interaction to adaptively handle different degradation types, while OneRestore(Guo et al., 2024a) aligns text-based degradation categories with visual features as prompts. IDR(Zhang et al., 2023a) uses meta-learning to decompose degradations, enabling generalization across diverse scenarios. MoFME(Zhang et al., 2024b), on the other hand, employs an uncertainty-aware router to dynamically assign Degradation-specific features to expert modules, allowing it to specialize in multiple tasks. Similarly, AdaIR(Cui et al., 2025) achieves adaptive restoration by decoupling degradations in the frequency domain. Additionally, methods like DA-CLIP(Luo et al., 2023a) leverage the pre-trained vision-language model (VLM) for prompt extraction, but the significant computational overhead introduced by VLM limits their practical efficiency. However, these methods either limit the representation of degradation information to coarse-grained degradation categories or lack a clear mechanism for extracting degradation features, which often results in suboptimal performance. In contrast, we propose a novel All-in-One image restoration framework, DPMambaIR, which is capable of extracting fine-grained degradation information and incorporating a state-space modeling process, effectively addressing diverse degradation challenges in image restoration.
2.2. State Space Model
State Space Model (SSM), originally introduced in the 1960s for control systems(Kalman, 1960), has recently gained significant attention in computer vision due to its ability to model long-range dependencies with linear computational complexity. The first effort to incorporate SSM into vision tasks was initiated by Vision Mamba(Zhu et al., 2024), which proposed the Vision State Space Module (VSSM) to extend SSM to visual applications. Vision Mamba demonstrated superior performance compared to Vision Transformers (ViTs)(Dosovitskiy et al., 2020), while maintaining lower model complexity, thereby laying a solid foundation for subsequent advancements in the field of computer vision. The success of VSSM has inspired extensive research into the application of SSM across various vision tasks, such as object detection(Zhang et al., 2024a; Yang et al., 2025; Liu et al., 2024; Li et al., 2024c; Zhang et al., 2023b, c), semantic segmentation(Xiao et al., 2024; Huang et al., 2023; Zhang et al., 2024c), and image classification(Zhang et al., 2024d; Li et al., 2024b; Yao et al., 2024; jc5sheng2024dualmamba). Among these, MambaIR(Guo et al., 2024c) is the first approach to apply SSM to image restoration. By leveraging SSM, MambaIR achieves enhanced computational efficiency and improved global feature modeling capabilities, resulting in significant improvements in perceptual quality for restoration tasks. Furthermore, other SSM-based approaches have also emerged for various image restoration tasks(Hu and Yin, 2024; Li et al., 2024d; Guo et al., 2024b; He et al., 2024c; Sun et al., 2024b; Li et al., 2024a; Wu et al., 2024c; Xia et al., 2024; Peng et al., 2025a; Di et al., 2024; Peng et al., 2025b; Pan et al., 2025b; Qi et al., 2025; Sun et al., 2024a; Wu et al., 2024b; He et al., 2024a; Du et al., 2024; He et al., 2024d; Jiang et al., 2024; Wang et al., 2024a; Li et al., 2023; Conde et al., 2024; Ren et al., 2024; Yan et al., 2023), showcasing the strong potential and promising opportunities of SSM in this domain. However, while SSM excels at modeling long-range interactions, they inherently lack a degradation-aware mechanism, which is crucial for tasks that involve handling diverse degradations. This limitation makes it difficult to balance differences between multiple tasks during modeling. This limitation hinders the performance of the model in All-in-One image restoration tasks.
3. Method
3.1. Problem Definition
All-in-One Image Restoration aims to address a wide range of image degradation types within a unified framework. Specifically, given a degraded image $\boldsymbol{I}_{D}\in\mathbb{R}^{3\times H\times W}$, where $D$ denotes the unknown degradation process, our goal is to restore the clean image $\boldsymbol{\hat{O}}\in\mathbb{R}^{3\times H\times W}$ using a unified model $\mathcal{F}$. This can be formulated as:
(1) $\boldsymbol{\hat{O}}=\mathcal{F}(\boldsymbol{I}_{D}).$
where $f$ is designed to generalize across various degradation types, such as noise, blur, compression artifacts, and weather effects, without explicit knowledge of $D$. The degradation process $D$ can be mathematically represented as:
(2) $\boldsymbol{I}_{D}=D(\boldsymbol{I}_{C})+\boldsymbol{\eta}.$
Here, $\boldsymbol{I}_{C}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $D$ represents the degradation operator (e.g., blur kernel, down-sampling, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise. The main challenge arises from the fact that $D$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
3.2. Traditional State Space Model
In Image Restoration, Convolutional Neural Networks (CNN)(LeCun et al., 2015) and Transformer(Vaswani et al., 2017) have been widely adopted due to their strong feature extraction and representation learning capabilities. CNN excels at capturing local patterns, making them suitable for tasks requiring high spatial resolution and fine details. However, their limited receptive field often hinders their ability to model long-range dependencies, which are critical for handling global degradations. On the other hand, Transformer leverages self-attention mechanisms to capture global context effectively. Despite their success, Transformer suffers from quadratic complexity with respect to the input size, leading to high computational costs and memory consumption, especially for high-resolution images. To address these limitations, researchers have turned to Mamba, which leverages State Space Model (SSM) to achieve linear complexity and long-range dependency modeling, making it well-suited for Image Restoration tasks.
Let’s briefly review the traditional State Space Model (SSM), which maps sequence input $x(t)$ to output $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$ and can be represented as a linear ordinary differential equation:
(3) $\dot{h}(t)$ $=Ah(t)+Bx(t),$
$y(t)$ $=Ch(t)+Dx(t).$
where $N$ is the state size, $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}$. Discretized using a zero-order hold as:
(4) $\overline{A}$ $=\exp(\Delta A),$
$\overline{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B.$
After the discretization, the discretized version of Eq. (3) with step size $\Delta$ can be rewritten as:
(5) $h_{i}$ $=\overline{A}h_{i-1}+\overline{B}x_{i},$
$y_{i}$ $=Ch_{i}+Dx_{i}.$
The advent of State Space Model has garnered significant attention for its ability to perform global modeling with linear complexity in image restoration. Existing Mamba-based methods(Zhu et al., 2024; Guo et al., 2024c, b) typically adopt specific scanning strategies to unfold 2D images into 1D sequences, enabling sequential causal modeling using Mamba. However, in All-in-One image restoration tasks, traditional SSM still faces two key limitations: (a) Limited Global Context Utilization: During the sequential causal modeling process, the $i$-th pixel can only access the previous $i-1$ pixels in the sequence, preventing effective utilization of global contextual information. To address this limitation, existing methods often employ parallel multi-directional scanning to expand the receptive field. However, this inevitably increases computational complexity. Furthermore, experiments in existing work(Guo et al., 2024c) reveal that parallel multi-directional scanning contributes minimally to performance improvements in low-level vision tasks. (b) Lack of Effective Degradation-Aware Mechanisms: Traditional SSM lack mechanisms to effectively perceive degradation information, making it difficult to balance differences between multiple tasks during modeling. This limitation hinders the performance of the model in All-in-One image restoration tasks.
3.3. DPMambaIR
We propose DPMambaIR, a degradation-aware framework for All-in-One image restoration. As illustrated in Figure 2, we first establish a Mamba-based image restoration baseline using an asymmetric encoder-decoder design within a U-shaped architecture. The restoration process begins with a $3\times 3$ convolutional layer, which extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded image. These features then pass through three encoding stages, each followed by downsampling. During each encoding stage, the resolution of the feature map is halved while the number of channels is doubled. Subsequently, the features pass through three decoding stages, where each stage involves upsampling and concatenating the output of the decoder with the corresponding output of the encoder to form the input for the next decoding stage. Finally, after a refinement stage, the clean image is restored through a shortcut connection that combines the input degraded image with the output of the decoder.
To address the challenges posed by unknown degradation types and levels in All-in-One image restoration tasks, we propose a Degradation Extractor pre-trained using degraded image reconstruction. This model compresses degradation type and severity information into a compact embedding, enabling the model to transition from blind to non-blind image restoration through finer-grained degradation representation. To further optimize the utilization of the extracted degradation embedding, we revisit the selective state-space modeling process. Specifically, we propose a Degradation-Aware Prompt State Space Model, which guides the image restoration process by compressing historical information and supplementing global degradation information.
3.4. Pretrained Degradation Extractor
Figure 3. (a) Degradation Modeling Process of Real-World Images, (b) The Architecture of Degradation Extractor.
Figure 4. The Architecture of (a) Degradation-aware Prompt Selective Scan (DPSS) and (b) High-frequency Enhancement Block (HEB), with Three Inputs $F_{1}$, $F_{2}$, and Degradation Embedding $D$.
To better extract degradation information from images, we revisit the degradation modeling process of real-world images, which spans from image capture and imaging to transmission, as illustrated in Figure  3 (a). This process typically involves four common types of degradations: (1) poor illumination conditions, (2) relative motion between the target objects and the camera, (3) artifacts introduced by transmission media in the optical path, and (4) compression during storage and transportation. We extend these four types of degradations to a more general framework and propose a unified degradation imaging model. The specific formulation is defined as follows:
(6) $I(x)=Q\big{(}(\alpha(x)J(x)+\beta(x)\gamma(x))\otimes K\big{)}.$
where $J(x)$ represents the clean image, $\alpha(x)$ models the luminance adjustment, such as the effects of insufficient lighting in low-light or adverse weather conditions (e.g., rain or fog), and $\gamma(x)$ accounts for artifacts caused by occlusions from aerosols, raindrops, or snow particles in the transmission medium. $\beta(x)$ modulates $\gamma(x)$ based on varying spatial properties, $K$ represents the blur introduced by relative motion between the target and the camera, and $Q$ denotes the compression applied during storage and transmission.
Based on the aforementioned degradation modeling process, we design a universal degradation extractor, as illustrated in Figure 3 (b). Specifically, we introduce central difference convolution(CDC)(Wang et al., 2023a; Peng et al., 2024a, b) to extract edge and texture degradation cues related to illumination and blur. In parallel, we employ three groups of CDC with different scales to capture artifacts of varying sizes, such as visual artifacts and snowflakes. Additionally, we integrate a traditional convolution branch to learn brightness-related representations.
Furthermore, to train the degradation extractor with fine-grained degradation encoding capability, we propose a pretraining method for the degradation extractor based on degraded image reconstruction. This method leverages a hybrid dataset consisting of paired data with diverse types of degradation. The core of this process is training the degradation extractor $\mathcal{E}$ to extract fine-grained degradation information from the degraded image $\boldsymbol{I}_{D}$ and distill it into a compact degradation embedding $D$ of size $1\times 512$, formulated as $D=\mathcal{E}(\boldsymbol{I}_{D})$. Subsequently, the reconstruction unit $\mathcal{R}$ utilizes $D$ and the clean image $\boldsymbol{O}$ to reconstruct the degraded image $\boldsymbol{I}_{D}$, formulated as $\boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{O},D)$.
We further introduce a dual-component loss function $L$, which combines the $L_{1}$ norm and the Learned Perceptual Image Patch Similarity (LPIPS) metric. The combined loss function is expressed as $L(\boldsymbol{I}_{D},\boldsymbol{\hat{I}}_{D})=L_{1}+L_{LPIPS}$, striking a careful balance between reconstruction accuracy and perceptual quality.
3.5. Degradation-aware Prompt State Space Model
To address these limitation mentioned in 3.2, we propose a novel Degradation-aware Prompt State Space Model to fully unlock the potential of SSM in All-in-One image restoration. Specifically, we propose a novel Degradation-aware Prompt Selective Scan, which incorporates degradation information into the State Space Model. This is achieved by leveraging a Degradation Embedding extracted from a pre-trained Degradation Extractor to provide supplementary degradation-related contextual information that the traditional SSM cannot capture. In addition, we modulate the parameter matrices in the state-space equations to establish a degradation-aware mechanism. This mechanism enables SSM to adapt to the imbalances caused by the inherent differences among sub-tasks in All-in-One image restoration. To further mitigate the issue of high-frequency information loss, which arises from balancing the differences among sub-tasks during training, we design a simple yet effective High-frequency Enhancement Block (HEB). This component ensures the preservation of high-frequency details and significantly enhances the overall performance of the model.
Degradation-aware Prompt Selective Scan. We revisit the traditional SSM, whose parameter settings are shown in Equation (3). Matrices $A$, $B$, $C$, and $D$ control the SSM output. The state transition matrix $A$ compresses historical information, while the input matrix $B$ maps input signals to the state space, determining their influence on historical states. The output matrix $C$ maps historical states to the observable space, reflecting their impact on the output. Matrix $D$ provides a direct pathway from input to output, similar to a residual connection. In traditional SSM, these matrices and $\Delta$ are derived from:
(7) $A=init(),D=init(),$
$B=Linear(x),C=Linear(x),\Delta=Linear(x).$
where $init()$ is an initialization method that does not depend on the input $x$. This way may lead to two key issues: (a) Traditional State Space Model struggles to effectively leverage global contextual information. (b) The lack of a proper degradation-aware mechanism makes it challenging to balance the differences among tasks in multi-task learning.
To address these issues, we first utilize a pre-trained degradation extractor $\mathcal{E}$ to extract the degradation embedding $D$ from the input image $I_{D}$. Subsequently, $D$ is used to modulate $\Delta$, the input matrix $B$, and the output matrix $C$ of the SSM. This introduces a degradation-aware mechanism in three aspects: historical information compression, input information processing, and global context awareness, as illustrated in Figure 4 (a). The specific method is formulated as follows:
(8) $D$ $=\mathcal{E}(I_{D}),$
$\alpha_{\Delta}$ $=\mathcal{M}_{\Delta}(D),\,\alpha_{B}=\mathcal{M}_{B}(D),\,\alpha_{C}=\mathcal{M}_{C}(D),$
$\Delta_{dp}$ $=\alpha_{\Delta}\cdot\Delta,\,B_{dp}=\alpha_{B}\cdot B,\,C_{dp}=\alpha_{C}\cdot C.$
Here, $\mathcal{M}_{\Delta}$, $\mathcal{M}_{B}$, and $\mathcal{M}_{C}$ are combinations of linear layers used to map the degradation embedding $\mathbf{e}$ into feature spaces, forming modulation vectors $\alpha_{\Delta}$, $\alpha_{B}$, and $\alpha_{C}$ for $\Delta$, $B$, and $C$, respectively. $\Delta_{dp}$, $B_{dp}$, and $C_{dp}$ are the newly modulated SSM parameters. Finally, the state-space function formulation of the Degradation-aware State Space Model can be expressed as follows:
(9) $\overline{A}_{dp}$ $=\exp(\Delta_{dp}A),$
$\overline{B}_{dp}$ $=(\Delta_{dp}A)^{-1}(\exp(A)-I)\cdot\Delta_{dp}B_{dp},$
$h_{i}$ $=\overline{A}_{dp_{i}}h_{i-1}+\overline{B}_{dp_{i}}x_{i},$
$y_{i}$ $=C_{dp}h_{i}+Dx_{i}.$
Through this approach, the transformation matrix $A$ and input matrix $B$ gain degradation-aware capabilities, integrating degradation cues of various types and levels during historical information compression and input signal processing, making the model more task-friendly. The output matrix $C$ controls how historical information affects the output. In traditional SSM, the $i$-th pixel can only observe the $i-1$ preceding pixels, leading to a limited global receptive field. While existing methods address this with parallel multi-directional scanning, this increases computational costs and offers limited performance gains in low-level vision tasks (Guo et al., 2024c). Instead, we retain single-directional scanning and enhance global degradation cues via $C$, achieving efficient global awareness with lower computational overhead.
High-frequency Enhancement Block.
Figure 5. We compare frequency-domain differences between clear and restored images under three configurations. The All-in-One setting shows a notable high-frequency gap compared to the Degradation-specific setting, which is significantly reduced by adding the High-frequency Enhancement Block (HEB), as shown by the blue line.
In All-in-One image restoration tasks, models tend to prioritize low-frequency restoration over high-frequency details, such as edges and textures, leading to their loss, as shown in Figure 5. To address this, we propose an efficient high-frequency enhancement block with minimal computational overhead. For input features $F$, we first extract channel-wise low-frequency features using global average pooling, then subtract them from $F$ to obtain high-frequency features, as shown in Figure 4(b). Finally, the high-frequency features are added back to enhance $F$, formulated as:
(10) ${F}_{enhanced}={F}+\alpha\cdot\big{(}{F}-\mathcal{G}({F})\big{)}.$
Here, $\alpha$ represents a learnable parameter that controls the weight of high-frequency information during the enhancement process. $\mathcal{G}$ represents the global average pooling operation. ${F}_{enhanced}$ denotes the features after high-frequency enhancement.
3.6. Loss Function
We employ the commonly used L1 loss $\mathcal{L}_{1}$ and L2 loss $\mathcal{L}_{2}$ in low-level vision tasks. Additionally, following previous works(Zamir et al., 2022; Cui et al., 2023), we incorporate a frequency-domain loss $\mathcal{L}_{fft}$ for training. The total loss is defined as:
(11) $\mathcal{L}_{total}=\lambda_{1}\cdot\mathcal{L}_{1}(O,\hat{O})+\lambda_{2}\cdot\mathcal{L}_{2}(O,\hat{O})+\lambda_{3}\cdot\mathcal{L}_{fft}(O,\hat{O}).$
where $O$ and $\hat{O}$ denote the model output and the ground truth, respectively. The parameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are balancing factors, which we set to 1, 0.5 and 0.001, respectively.
4. Experiment And Analysis
We evaluate our method on both All-in-one image restoration tasks and Degradation-specific image restoration tasks. In the All-in-One Image Restoration setting, we train a single model on a dataset containing seven types of degradation. For Degradation-specific image restoration, we train separately on datasets corresponding to each type of degradation.
4.1. Experimental Settings
Training Details. Building on previous works(Zamir et al., 2022; Potlapalli et al., 2023; Guo et al., 2024c), we develop a Mamba-based baseline for image restoration. Our method is trained using the AdamW(Kingma and Ba, 2014) optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $1e^{-4}$ for 300,000 iterations. The initial learning rate of $3e^{-4}$ is gradually reduced to $1e^{-6}$ through cosine annealing. Following(Zamir et al., 2022), we employ a progressive training strategy, cropping image patches of size $192\times 192$. Horizontal and vertical flips are applied for data augmentation. Our experiments are conducted on a single NVIDIA RTX A100 GPU and implemented using the PyTorch platform.
Datasets. For All-in-One image restoration, we collect a large-scale mixed dataset that includes seven classic and popular degradation types: fog, rain, snow, low light, blur, JPEG compression artifacts, and noise. Additionally, we evaluate performance on Degradation-specific tasks, using the LOL dataset(Wei et al., 2018) for low-light image enhancement, the GoPro dataset(Nah et al., 2017a) for image deblurring, the Rain100H dataset(Yang et al., 2017) for image deraining, and the RESIDE6K dataset(Qin et al., 2020) for image dehazing.
Table 1. Quantitative Comparison on Seven Datasets for All-in-One Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Method Lowlight Snow Rain Fog Jpeg Noise Blur Average PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM MPRNet 19.71 0.659 30.94 0.938 26.27 0.912 25.63 0.947 30.51 0.938 28.03 0.897 26.07 0.869 26.74 0.880 MIRNet 19.81 0.668 31.04 0.941 25.77 0.907 22.26 0.882 30.54 0.939 28.14 0.900 26.32 0.873 26.27 0.873 NAFNet 19.49 0.650 30.33 0.935 26.04 0.906 25.63 0.946 30.20 0.936 27.80 0.895 26.13 0.870 26.52 0.877 Restormer 19.99 0.670 31.67 0.944 26.91 0.919 26.18 0.953 30.56 0.939 28.17 0.890 26.62 0.880 27.16 0.886 PromptIR 19.90 0.669 31.52 0.943 26.80 0.918 26.13 0.945 30.57 0.939 28.18 0.901 26.36 0.874 27.07 0.884 IDR 19.68 0.663 31.16 0.941 26.39 0.914 26.11 0.948 30.51 0.939 28.10 0.899 26.29 0.873 26.89 0.883 OneRestore 19.77 0.657 29.95 0.931 25.37 0.900 27.74 0.962 30.10 0.934 27.65 0.890 25.90 0.866 26.64 0.877 AdaIR 19.99 0.672 31.94 0.923 27.01 0.922 26.61 0.952 30.53 0.939 28.19 0.900 26.61 0.879 27.26 0.887 DPMambaIR 20.04 0.680 32.36 0.949 27.07 0.926 28.13 0.963 30.57 0.939 28.25 0.902 27.43 0.894 27.69 0.893
Table 2. Quantitative Comparison on Deraining, Low-light Image Enhancement(LLIE), Deblurring and Dehazing for Degradation-Specific Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Deraining PSNR SSIM JORDER 26.25 0.835 PReNet 29.46 0.899 MPRNet 30.41 0.891 MAXIM 30.81 0.903 Restormer 31.46 0.904 DPMambaIR 32.30 0.967
LLIE PSNR SSIM EnlightenGAN 17.61 0.653 MIRNet 24.14 0.830 URetinex-Net 19.84 0.824 MAXIM 23.43 0.863 IAT 23.38 0.809 DPMambaIR 24.20 0.852
Deblurring PSNR SSIM DeepDeBlur 29.08 0.913 DeBlurGAN 28.70 0.858 DeBlurGANV2 29.55 0.934 MT-RNN 31.15 0.945 IR-SDE 30.70 0.901 DPMambaIR 31.42 0.948
Dehazing PSNR SSIM GCANet 26.59 0.935 GridDehazeNet 25.86 0.944 DeHazeFormer 30.29 0.964 MAXIM 29.12 0.932 DA-CLIP 30.16 0.936 DPMambaIR 30.89 0.977
Figure 6. Qualitative comparison on Deblurring, Deraining and Dehazing for All-in-One Image Restoration Task.
Evaluation Metrics. Following previous works(Zamir et al., 2022; Guo et al., 2024c), we use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively compare different methods on the RGB channels.
Comparisons with State-of-the-art Methods. For All-in-One image restoration, we compare our method against eight approaches: MIRNet (Zamir et al., 2020), NAFNet (Chen et al., 2022), MPRNet (Zamir et al., 2021), Restormer (Zamir et al., 2022), PromptIR (Potlapalli et al., 2023), IDR (Zhang et al., 2023a), OneRestore (Guo et al., 2024a), and AdaIR (Cui et al., 2025). Among these, MIRNet, NAFNet, MPRNet, and Restormer are general-purpose image restoration models, while the latter four methods are specifically designed for All-in-One image restoration. Specifically, PromptIR employs visual prompts to guide restoration, OneRestore leverages a pre-trained degradation category encoder with cross-attention mechanisms, and AdaIR utilizes frequency domain information for adaptive restoration. For Degradation-specific image restoration tasks, we compare our method with different state-of-the-art approaches tailored to each domain. The details are as follows: (1) For deraining, we compare against JORDER (Yang et al., 2019), PReNet (Ren et al., 2019), MPRNet (Zamir et al., 2021), MAXIM (Tu et al., 2022), and Restormer (Zamir et al., 2022). (2) For low-light image enhancement, we compare against EnlightenGAN (Jiang et al., 2021), MIRNet (Zamir et al., 2020), Uretinex-Net (Wu et al., 2022), MAXIM (Tu et al., 2022), and IAT (Cui et al., 2022). (3) For deblurring, we compare against DeepDeblur (Nah et al., 2017b), DeblurGAN (Kupyn et al., 2017), DeblurGANv2 (Kupyn et al., 2019), MT-RNN (Park et al., 2020), and IR-SDE (Luo et al., 2023b). (4) For dehazing, we compare against GCANet (Chen et al., 2018), GridDehazeNet (Liu et al., 2019), DeHazeFormer (Song et al., 2023), MAXIM (Tu et al., 2022), and DA-CLIP (Luo et al., 2023a).
4.2. Quantitative Results
All-in-one Image Restoration. We compare our method against five state-of-the-art image restoration approaches, including both CNN-based and transformer-based methods. As shown in Table 1, for the All-in-One image restoration task, our method achieves the best PSNR and SSIM on the aforementioned mixed datasets compared to existing methods. Specifically, our method achieves a PSNR of 27.69dB and an SSIM of 0.893, outperforming AdaIR by 0.43dB and 0.011, respectively. We achieve the best performance across all seven sub-tasks. Experimental results indicate that multi-task learning often encounters challenges in balancing performance across different tasks. For instance, AdaIR demonstrates strong performance in deraining but performs poorly in dehazing. On the other hand, OneRestore excels in dehazing but delivers suboptimal results in tasks such as deraining and deblurring. In contrast, our method effectively tackles the balance issue in multi-task learning, achieving consistently optimal performance across all sub-tasks. These findings highlight the effectiveness and superiority of our proposed method.
Degradation-Specific Image Restoration. We evaluate our method on four Degradation-specific tasks, consistently achieving superior performance across multiple datasets compared to state-of-the-art CNN-based, transformer-based, and diffusion-based approaches (Table 2). On Rain100H (Yang et al., 2017), our method achieves a PSNR of 32.30dB and SSIM of 0.967, surpassing Restormer by 0.84dB and 0.063, respectively. On the LOL dataset (Wei et al., 2018), it achieves 24.20dB PSNR and 0.852 SSIM, outperforming MIRNet by 0.06dB and 0.022. For deblurring on GoPro (Nah et al., 2017a), it attains 31.42dB PSNR and 0.948 SSIM, exceeding MT-RNN by 0.27dB and 0.003. On RESIDE6K (Qin et al., 2020), it achieves 30.89dB PSNR and 0.973 SSIM, outperforming DehazeFormer by 0.60dB and 0.009. These results highlight the versatility and effectiveness of our method in addressing diverse degradation restoration tasks.
4.3. Qualitative Results
Figure 6 presents the visual results of deblurring, deraining, and dehazing under the All-in-One Image Restoration task setting. Our method demonstrates superior visual performance compared to other approaches. Specifically, in deblurring, our method restores sharper edge details; in deraining, it more effectively removes rain streaks while recovering background information; and in dehazing, it focuses on regions often overlooked by other methods, resulting in more comprehensive dehazing performance. Additional visual comparisons under both the All-in-One and Degradation-specific settings are provided in the Appendix.
4.4. Ablation Study
Table 3 presents the results of our ablation study on DPMambaIR for the All-in-One image restoration task, illustrating the impact of removing the degradation prompt, different modulation modes, and the High-frequency Enhancement Block (HEB) from the model. For the baseline model without the Degradation-Aware Prompt Selective Scan and HEB, the PSNR and SSIM are only 26.96 dB and 0.884, respectively. Introducing degradation-awareness via prompts applied to Matrices $A$ and $B$ improves PSNR and SSIM by 0.64 dB and 0.008, respectively, by enabling degradation-aware mechanisms in historical information compression and input feature processing. Adding Prompt $C$, which supplements global degradation information, further increases PSNR by 0.07 dB. The inclusion of HEB significantly boosts performance, preserving high-frequency details. The complete model achieves 27.69 dB PSNR and 0.893 SSIM, demonstrating the importance of all components for optimal performance. Table 4 presents the results for degradation extraction and utilization methods. For degradation extraction, we compare three approaches: explicit category labels, a pre-trained degradation extractor from OneRestore (Guo et al., 2024a), and our fine-grained degradation representation via image reconstruction. The OneRestore extractor, designed for classification tasks, performs worse than explicit category labels. In contrast, our method achieves a 0.11 dB PSNR improvement over non-blind restoration with category labels by leveraging fine-grained degradation representations. For degradation utilization, we compare concatenation and attention, with our method outperforming both, demonstrating its effectiveness and superiority in extracting and utilizing degradation information.
For the ablation study on global information supplementation, we compare our method with Mamba’s dual-sequence scanning scheme. As shown in Table 5, our approach not only achieves superior performance compared to the dual-sequence scanning method but also demonstrates a significant reduction in both the Parameters and FLOPs.
Table 3. Ablation Studies on our proposed core Modules.
[…]
Table 4. Ablation Study on Different Degradation Extraction and Utilization Methods.
Comparison Projects Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Ours 27.69 0.893
Table 5. Ablation Study on Different Methods for Global Information Supplementation.
Method PSNR SSIM Params(M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt Matrix C 27.69 0.893 29.35 153.3
5. Conclusion
In this paper, we proposed DPMambaIR, a novel All-in-One framework for image restoration that balances high restoration quality and computational efficiency across diverse degradation types. DPMambaIR introduces a Degradation-Aware Prompt State Space Model (DP-SSM), which dynamically integrates fine-grained degradation features into the state-space modeling process. This design enables adaptive handling of complex degradation scenarios while maintaining global degradation awareness. Furthermore, the High-Frequency Enhancement Block (HEB) reduces high-frequency detail loss caused by task competition, ensuring superior edge and texture restoration. This study underscores the importance of fine-grained degradation modeling and dynamic feature modulation in advancing All-in-One image restoration frameworks. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, establishing it as a promising approach for unified image restoration.

### v2

@@
 DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model
-Abstract.
-All-in-One image restoration aims to address multiple image degradation problems using a single model, significantly reducing training costs and deployment complexity compared to traditional methods that design dedicated models for each degradation type.
+Abstract
+All-in-One image restoration aims to address multiple image degradation problems using a single model, offering a more practical and versatile solution compared to designing dedicated models for each degradation type.
 Existing approaches typically rely on Degradation-specific models or coarse-grained degradation prompts to guide image restoration.
 However, they lack fine-grained modeling of degradation information and face limitations in balancing multi-task conflicts.
-To overcome these limitations, we propose DPMambaIR, a novel All-in-One image restoration framework.
-By integrating a Degradation-Aware Prompt State Space Model (DP-SSM) and a High-Frequency Enhancement Block (HEB), DPMambaIR enables fine-grained modeling of complex degradation information and efficient global integration, while mitigating the loss of high-frequency details caused by task competition.
-Specifically, the DP-SSM utilizes a pre-trained degradation extractor to capture fine-grained degradation features and dynamically incorporates them into the state space modeling process, enhancing the model’s adaptability to diverse degradation types.
-Concurrently, the HEB supplements high-frequency information, effectively addressing the loss of critical details, such as edges and textures, in multi-task image restoration scenarios.
+To overcome these limitations, we propose DPMambaIR, a novel All-in-One image restoration framework that introduces a fine-grained degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM).
+The DP-SSM leverages the fine-grained degradation features captured by the extractor as dynamic prompts, which are then incorporated into the state space modeling process.
+This enhances the model’s adaptability to diverse degradation types, while a complementary High-Frequency Enhancement Block (HEB) recovers local high-frequency details.
 Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, with 27.69dB and 0.893 in PSNR and SSIM, respectively.
 These results highlight the potential and superiority of DPMambaIR as a unified solution for All-in-One image restoration.
-All-in-One Image Restoration, State Space Model, Prompt Learning
-1. Introduction
+I Introduction
 Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs.
-Recent advances in deep learning have significantly improved performance in Degradation-specific restoration problems, such as image denoising (Kim et al., 2024; Zeng et al., 2024; Ryou et al., 2024), deblurring (Chen et al., 2024b; Wu et al., 2024a; Mao et al., 2024), deraining(Chen et al., 2024a, 2023; Ba et al., 2022; Peng et al., 2020, 2021; He et al., 2024b), dehazing(Zhang et al., 2024e; Wang et al., 2024c; Zheng et al., 2023; Wu et al., 2023) and low-light enhancement(Ma et al., 2022; Cai et al., 2023; Hai et al., 2023; Wang et al., 2022, 2023b).
+Recent advances in deep learning have significantly improved performance in Degradation-specific restoration tasks, such as image denoising [1, 2], deblurring [3], deraining[4], dehazing[5, 6] and low-light enhancement[7].
 However, traditional methods typically rely on Degradation-specific models explicitly tailored to individual degradation types.
 These approaches often require prior knowledge of the degradation type, which limits their practicality in real-world scenarios, such as autonomous driving and nighttime surveillance, where diverse and unknown degradations frequently coexist.
-Figure 1. Comparison with SOTA All-in-One Image Restoration Methods on mixed Dataset. Our DPMambaIR achieves consistently better performance.
+Figure 1: Comparison with SOTA All-in-One Image Restoration Methods on mixed Dataset. Our DPMambaIR achieves consistently better performance.
 To address this limitation, recent studies explore All-in-One image restoration frameworks that unify multiple restoration tasks into a single model.
 Existing methods can be broadly categorized into Mixture-of-Experts (MoE)-based and Prompt-based approaches.
-MoE-based methods, such as MoFME (Zhang et al., 2024b), dynamically route tasks to specialized experts using an uncertainty-aware router, enabling effective learning of diverse expert functions.
+MoE-based methods, such as MoFME [8], dynamically route tasks to specialized experts using an uncertainty-aware router, enabling effective learning of diverse expert functions.
 However, these methods suffer from high computational costs, increased inference complexity as the number of experts grows, and lack explicit degradation modeling, limiting adaptability to unseen or complex degradations.
-Prompt-based methods, like OneRestore (Guo et al., 2024a), guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms for flexible and adaptive strategies.
+Prompt-based methods, like OneRestore [9], guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms for flexible and adaptive strategies.
 While effective, these methods rely on coarse prompts that fail to capture nuanced degradation characteristics and treat prompts as static inputs, limiting their ability to adapt to varying degradation intensities or specific requirements, leading to suboptimal results.
-While these approaches reduce computational cost and deployment complexity by leveraging feature sharing and multi-task learning, two key challenges remain unresolved.
-First, existing models lack fine-grained degradation awareness and often rely on coarse degradation labels to guide the restoration process.
+While these methods have demonstrated success in learning multiple tasks for All-in-One image restoration, two key challenges remain unresolved.
+First, existing models lack a degradation extractor capable of capturing fine-grained degradation features and often rely on coarse degradation labels to guide the restoration process.
 For example, images with varying degrees of haze are typically assigned a single haze label, failing to capture the nuanced variations in degradation levels.
@@
 Second, current methods struggle to balance the competing requirements of varying restoration tasks.
-For instance, training low-light enhancement, which focuses on restoring low-frequency contrast information, alongside deblurring, which prioritizes the recovery of fine local details, often results in task competition.
+For instance, training low-light enhancement, which focuses on restoring low-frequency contrast information, alongside deblurring, which prioritizes the recovery of fine local details.
 This inherent task competition makes it challenging to optimize a single network for all restoration tasks simultaneously and remains a fundamental challenge for All-in-One image restoration frameworks.
-To address these challenges, we propose a novel All-in-One image restoration framework, DPMambaIR.
-The framework is built upon a Degradation-aware Prompt State Space Model (DP-SSM) and a lightweight High-frequency Enhancement Block (HEB).
-Specifically, DP-SSM employs a pre-trained degradation extractor to extract degradation embeddings, which are used to modulate the state transition, input, and output matrices of the state space equations.
-This enables fine-grained degradation modeling and efficient global degradation integration.
-Meanwhile, HEB mitigates the loss of high-frequency details caused by task competition among different restoration tasks.
+To address these challenges, we propose DPMambaIR, a novel All-in-One image restoration framework built upon a regression-based degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM).
+Unlike prior approaches that use static or class-based prompts, our method employs a degradation extractor to generate prompt vectors that encode both semantic categories and degradation distribution characteristics.
+These prompts are then used to dynamically modulate the state transition, input, and output matrices of a state space model, enabling fine-grained degradation modeling and adaptive coordination across frequency bands.
+This mechanism allows our model to better manage task competition and enhance restoration quality across diverse degradations.
+Moreover, while State Space Model such as Mamba exhibit strong global modeling capabilities, they are inherently biased toward low-frequency restoration [10].
+To better recover local high-frequency details, we introduce a lightweight High-Frequency Enhancement Block (HEB) as a supplementary component, which delivers performance gains while incurring only negligible computational overhead.
 Extensive experiments on a mixed dataset containing seven degradation types demonstrate the superiority of DPMambaIR.
-As shown in Figure 1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore.
+As shown in Fig.
+1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore.
 Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, showcasing its potential as an effective and robust All-in-One image restoration solution.
-The contributions of this paper can be summarized:
+The contributions of this paper are summarized as follows:
   • $\bullet$
@@
   • $\bullet$
-    We propose a Degradation-aware Prompt State Space Model (DP-SSM) that incorporates fine-grained degradation features into the state space modeling process, enabling both global degradation modeling and fine-grained degradation awareness to effectively handle diverse degradation types within a unified framework.
+    We design a Degradation-aware Prompt State Space Model (DP-SSM), which enhances the long-range modeling capability of Mamba by integrating degradation priors at multiple structural levels.
   • $\bullet$
@@
 The results validate the effectiveness and superiority of the proposed approach.
-Figure 2.
-The overall framework of DPMambaIR consists of a shallow feature extractor and multiple DPSS blocks, with an upsampler implemented by pixel shuffle and a downsampler by pixel unshuffle.
-Additionally, a degradation extractor is included to extract degradation embeddings, which are then inserted into the DPSS blocks.
-2. Related Work
-2.1. All-in-One Image Restoration
+II Related Work
+II-A All-in-One Image Restoration
 Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs.
-Traditional methods, such as Dark Channel Prior (He et al., 2010) and Color Line Prior (Fattal, 2014), rely on handcrafted priors to model degradation but struggle with complex scenarios due to predefined assumptions.
-The rise of deep learning (LeCun et al., 2015) has significantly advanced tasks like such as image super-resolution(Peng et al., 2024c, d, 2025c; Wang et al., 2024b; Saharia et al., 2022; Gao et al., 2023), image denoising (Kim et al., 2024; Zeng et al., 2024; Ryou et al., 2024), deblurring (Chen et al., 2024b; Wu et al., 2024a; Mao et al., 2024), deraining(Chen et al., 2024a; Pan et al., 2023; Chen et al., 2023; Ba et al., 2022; Peng et al., 2020, 2021; He et al., 2024b), dehazing(Zhang et al., 2024e; Wang et al., 2024c; Zheng et al., 2023; Wu et al., 2023), desnowing(Guo et al., 2025; Chen et al., 2021; Zhang et al., 2021) and low-light enhancement(Ma et al., 2022; Cai et al., 2023; Hai et al., 2023; Wang et al., 2022, 2023b; Ma et al., 2024).
-These approaches, leveraging convolutional neural networks (CNN)-based (LeCun et al., 1989), transformer-based (Vaswani et al., 2017; Pan et al., 2025a, 2024), or diffusion-based (Ho et al., 2020) architectures, enable powerful feature learning for image restoration.
+Traditional methods, such as Dark Channel Prior [11] and Color Line Prior [12], rely on handcrafted priors to model degradation but struggle with complex scenarios due to predefined assumptions.
+The rise of deep learning [13] has significantly advanced tasks like such as image super-resolution[14, 15, 16], image denoising [1, 2], deblurring [3], deraining[4], dehazing[5], desnowing[17] and low-light enhancement[7].
+These approaches, leveraging convolutional neural networks (CNN)-based [18], transformer-based [19], or diffusion-based [20] architectures, enable powerful feature learning for image restoration.
 However, most existing works focus on addressing a single type of degradation, which limits their applicability in real-world scenarios where images often suffer from diverse and unknown types of degradation.
 Recent advancements in All-in-One image restoration frameworks focus on enhancing flexibility and efficiency by leveraging various multi-task learning techniques such as Mixture-of-Experts (MoE) and text-based or vision-based prompts to handle diverse restoration tasks.
-For example, AirNet(Li et al., 2022) employs contrastive learning to extract degradation representations for guiding restoration.
-PromptIR(Potlapalli et al., 2023) integrates end-to-end prompt generation and interaction to adaptively handle different degradation types, while OneRestore(Guo et al., 2024a) aligns text-based degradation categories with visual features as prompts.
-IDR(Zhang et al., 2023a) uses meta-learning to decompose degradations, enabling generalization across diverse scenarios.
-MoFME(Zhang et al., 2024b), on the other hand, employs an uncertainty-aware router to dynamically assign Degradation-specific features to expert modules, allowing it to specialize in multiple tasks.
-Similarly, AdaIR(Cui et al., 2025) achieves adaptive restoration by decoupling degradations in the frequency domain.
-Additionally, methods like DA-CLIP(Luo et al., 2023a) leverage the pre-trained vision-language model (VLM) for prompt extraction, but the significant computational overhead introduced by VLM limits their practical efficiency.
-However, these methods either limit the representation of degradation information to coarse-grained degradation categories or lack a clear mechanism for extracting degradation features, which often results in suboptimal performance.
-In contrast, we propose a novel All-in-One image restoration framework, DPMambaIR, which is capable of extracting fine-grained degradation information and incorporating a state-space modeling process, effectively addressing diverse degradation challenges in image restoration.
-2.2. State Space Model
-State Space Model (SSM), originally introduced in the 1960s for control systems(Kalman, 1960), has recently gained significant attention in computer vision due to its ability to model long-range dependencies with linear computational complexity.
-The first effort to incorporate SSM into vision tasks was initiated by Vision Mamba(Zhu et al., 2024), which proposed the Vision State Space Module (VSSM) to extend SSM to visual applications.
-Vision Mamba demonstrated superior performance compared to Vision Transformers (ViTs)(Dosovitskiy et al., 2020), while maintaining lower model complexity, thereby laying a solid foundation for subsequent advancements in the field of computer vision.
-The success of VSSM has inspired extensive research into the application of SSM across various vision tasks, such as object detection(Zhang et al., 2024a; Yang et al., 2025; Liu et al., 2024; Li et al., 2024c; Zhang et al., 2023b, c), semantic segmentation(Xiao et al., 2024; Huang et al., 2023; Zhang et al., 2024c), and image classification(Zhang et al., 2024d; Li et al., 2024b; Yao et al., 2024; jc5sheng2024dualmamba).
-Among these, MambaIR(Guo et al., 2024c) is the first approach to apply SSM to image restoration.
-By leveraging SSM, MambaIR achieves enhanced computational efficiency and improved global feature modeling capabilities, resulting in significant improvements in perceptual quality for restoration tasks.
-Furthermore, other SSM-based approaches have also emerged for various image restoration tasks(Hu and Yin, 2024; Li et al., 2024d; Guo et al., 2024b; He et al., 2024c; Sun et al., 2024b; Li et al., 2024a; Wu et al., 2024c; Xia et al., 2024; Peng et al., 2025a; Di et al., 2024; Peng et al., 2025b; Pan et al., 2025b; Qi et al., 2025; Sun et al., 2024a; Wu et al., 2024b; He et al., 2024a; Du et al., 2024; He et al., 2024d; Jiang et al., 2024; Wang et al., 2024a; Li et al., 2023; Conde et al., 2024; Ren et al., 2024; Yan et al., 2023), showcasing the strong potential and promising opportunities of SSM in this domain.
-However, while SSM excels at modeling long-range interactions, they inherently lack a degradation-aware mechanism, which is crucial for tasks that involve handling diverse degradations.
-This limitation makes it difficult to balance differences between multiple tasks during modeling.
-This limitation hinders the performance of the model in All-in-One image restoration tasks.
-3. Method
-3.1. Problem Definition
-All-in-One Image Restoration aims to address a wide range of image degradation types within a unified framework.
-Specifically, given a degraded image $\boldsymbol{I}_{D}\in\mathbb{R}^{3\times H\times W}$, where $D$ denotes the unknown degradation process, our goal is to restore the clean image $\boldsymbol{\hat{O}}\in\mathbb{R}^{3\times H\times W}$ using a unified model $\mathcal{F}$.
-This can be formulated as:
-(1) $\boldsymbol{\hat{O}}=\mathcal{F}(\boldsymbol{I}_{D}).$
-where $f$ is designed to generalize across various degradation types, such as noise, blur, compression artifacts, and weather effects, without explicit knowledge of $D$.
-The degradation process $D$ can be mathematically represented as:
-(2) $\boldsymbol{I}_{D}=D(\boldsymbol{I}_{C})+\boldsymbol{\eta}.$
-Here, $\boldsymbol{I}_{C}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $D$ represents the degradation operator (e.g., blur kernel, down-sampling, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise.
-The main challenge arises from the fact that $D$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
-3.2. Traditional State Space Model
-In Image Restoration, Convolutional Neural Networks (CNN)(LeCun et al., 2015) and Transformer(Vaswani et al., 2017) have been widely adopted due to their strong feature extraction and representation learning capabilities.
+For example, AirNet[21] employs contrastive learning to extract degradation representations for guiding restoration.
+PromptIR[22] integrates end-to-end prompt generation and interaction to adaptively handle different degradation types, while OneRestore[9] aligns text-based degradation categories with visual features as prompts.
+IDR[23] uses meta-learning to decompose degradations, enabling generalization across diverse scenarios.
+MoFME[8], on the other hand, employs an uncertainty-aware router to dynamically assign Degradation-specific features to expert modules, allowing it to specialize in multiple tasks.
+Similarly, AdaIR[24] achieves adaptive restoration by decoupling degradations in the frequency domain.
+Additionally, methods like DA-CLIP[25] leverage the pre-trained vision-language model (VLM) for prompt extraction, but the significant computational overhead introduced by VLM limits their practical efficiency.
+Despite their effectiveness, most prompt-based methods rely on coarse-grained class labels or static degradation tokens, which limits their ability to model continuous degradation distributions or adapt to image-specific degradation characteristics.
+In contrast, our method extracts fine-grained, regression-level degradation embeddings that encode both semantic and statistical degradation features.
+These embeddings are used to dynamically modulate the state-space parameters of the restoration backbone, enabling degradation-aware processing throughout the network.
+This facilitates better coordination across restoration tasks and reduces competition between frequency components.
+Figure 2: The overall framework of DPMambaIR consists of a shallow feature extractor and multiple DPSS blocks, with an upsampler implemented by pixel shuffle and a downsampler by pixel unshuffle.
+Additionally, a degradation extractor is included to extract degradation embeddings, which are then inserted into the DPSS blocks.
+II-B State Space Model based Image Restoration
+In Image Restoration, Convolutional Neural Networks (CNN) [13] and Transformer [19] have been widely adopted due to their strong feature extraction and representation learning capabilities.
 CNN excels at capturing local patterns, making them suitable for tasks requiring high spatial resolution and fine details.
@@
 On the other hand, Transformer leverages self-attention mechanisms to capture global context effectively.
-Despite their success, Transformer suffers from quadratic complexity with respect to the input size, leading to high computational costs and memory consumption, especially for high-resolution images.
-To address these limitations, researchers have turned to Mamba, which leverages State Space Model (SSM) to achieve linear complexity and long-range dependency modeling, making it well-suited for Image Restoration tasks.
-Let’s briefly review the traditional State Space Model (SSM), which maps sequence input $x(t)$ to output $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$ and can be represented as a linear ordinary differential equation:
-(3) $\dot{h}(t)$ $=Ah(t)+Bx(t),$
+Despite their success, Transformers suffer from quadratic complexity relative to the input size, making them computationally and memory-intensive.
+To address these limitations, researchers have turned to the State Space Model (SSM)[26], which provides an efficient framework for modeling long-range dependencies with linear computational complexity.
+Formally, SSM maps an input sequence $x(t)$ to an output sequence $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$, represented by a linear ordinary differential equation:
+$h^{\prime}(t)$ $=Ah(t)+Bx(t),$ (1)
 $y(t)$ $=Ch(t)+Dx(t).$
-where $N$ is the state size, $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}$. Discretized using a zero-order hold as:
-(4) $\overline{A}$ $=\exp(\Delta A),$
-$\overline{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B.$
-After the discretization, the discretized version of Eq. (3) with step size $\Delta$ can be rewritten as:
-(5) $h_{i}$ $=\overline{A}h_{i-1}+\overline{B}x_{i},$
+where $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}$.
+After discretization with step size $\Delta$, the dynamic equation can be expressed as:
+$h_{i}$ $=\overline{A}h_{i-1}+\overline{B}x_{i},$ (2)
 $y_{i}$ $=Ch_{i}+Dx_{i}.$
-The advent of State Space Model has garnered significant attention for its ability to perform global modeling with linear complexity in image restoration.
-Existing Mamba-based methods(Zhu et al., 2024; Guo et al., 2024c, b) typically adopt specific scanning strategies to unfold 2D images into 1D sequences, enabling sequential causal modeling using Mamba.
-However, in All-in-One image restoration tasks, traditional SSM still faces two key limitations: (a) Limited Global Context Utilization: During the sequential causal modeling process, the $i$-th pixel can only access the previous $i-1$ pixels in the sequence, preventing effective utilization of global contextual information.
-To address this limitation, existing methods often employ parallel multi-directional scanning to expand the receptive field.
-However, this inevitably increases computational complexity.
-Furthermore, experiments in existing work(Guo et al., 2024c) reveal that parallel multi-directional scanning contributes minimally to performance improvements in low-level vision tasks.
-(b) Lack of Effective Degradation-Aware Mechanisms: Traditional SSM lack mechanisms to effectively perceive degradation information, making it difficult to balance differences between multiple tasks during modeling.
-This limitation hinders the performance of the model in All-in-One image restoration tasks.
-3.3. DPMambaIR
+where $\overline{A}=\exp(\Delta A)$ and $\overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B$.
+The introduction of Vision Mamba[27] marked the first effort to incorporate SSM into vision tasks.
+By proposing the Vision State Space Module (VSSM), it extended SSM to visual applications and demonstrated superior performance over Vision Transformer (ViT)[28], while maintaining lower complexity.
+This success inspired extensive research into SSM across diverse tasks such as object detection[29], semantic segmentation[30], and image classification[31].
+In the context of Image Restoration, MambaIR[32] pioneered the integration of SSM, leveraging its linear complexity and long-range modeling capability to achieve significant improvements in both computational efficiency and perceptual quality.
+This was followed by several variants targeting specific restoration tasks[33, 34, 35], collectively showcasing the strong potential of SSM in this domain.
+Existing approaches typically unfold 2D images into 1D sequences and apply sequential causal modeling.
+However, this design still faces two limitations: (a) Limited Global Context Utilization: The $i$-th pixel can only access the previous i1i-1 pixels in the sequence, restricting the effective use of global context.
+Parallel multi-directional scanning has been explored to enlarge the receptive field, but it increases computational cost while providing limited gains in low-level vision tasks[32].
+(b) Lack of Effective Degradation-Aware Mechanisms: Current SSM-based methods primarily target single-task restoration.
+Their transition and output mappings are fixed or solely input-driven, without leveraging external degradation priors.
+As a result, they cannot dynamically adapt to diverse degradation types or levels, leading to suboptimal performance in All-in-One image restoration task.
+These limitations motivate the introduction of a degradation-aware modulation mechanism into the SSM pipeline.
+In this work, we propose a Degradation-aware Prompt State Space Model (DP-SSM), where degradation embeddings dynamically modulate the state-space transition and output matrices.
+This design enables a fine-grained, task-aware mechanism that balances global modeling with degradation-specific guidance, enhancing the flexibility and effectiveness of SSM for complex restoration scenarios.
+III Method
+III-A Problem Definition
+All-in-One Image Restoration aims to address a wide range of image degradation types within a unified framework.
+Specifically, given a degraded image $\boldsymbol{I}_{D}\in\mathbb{R}^{3\times H\times W}$, where $D$ denotes the degradation process, our goal is to restore the clean image $\boldsymbol{\hat{O}}\in\mathbb{R}^{3\times H\times W}$ using a unified model $\mathcal{F}$.
+This can be formulated as:
+$\boldsymbol{\hat{O}}=\mathcal{F}(\boldsymbol{I}_{D}).$ (3)
+where $\mathcal{F}$ is designed to generalize across various degradation types, such as noise, blur, compression artifacts, and weather effects, without explicit knowledge of $Deg$.
+The degradation process $Deg$ can be mathematically represented as:
+$\boldsymbol{I}_{D}=Deg(\boldsymbol{I}_{C})+\boldsymbol{\eta}.$ (4)
+Here, $\boldsymbol{I}_{C}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $Deg$ represents the degradation operator (e.g., blur kernel, brightness, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise.
+The main challenge arises from the fact that $Deg$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
+Unlike degradation-specific restoration, the All-in-One setting requires the model to generalize across a broad spectrum of degradation types without prior knowledge of degradation parameters.
+This introduces two main challenges: (a) how to represent fine-grained degradation information in a unified way; and (b) how to dynamically adapt the model behavior based on such information during inference.
+Figure 3: (a) Real-World Image Degradation Modeling Process.
+(b) Architecture details of the Large version degradation extractor and reconstructor.
+(c) Architecture details of the normal version degradation extractor and reconstructor.
+III-B DPMambaIR
 We propose DPMambaIR, a degradation-aware framework for All-in-One image restoration.
-As illustrated in Figure 2, we first establish a Mamba-based image restoration baseline using an asymmetric encoder-decoder design within a U-shaped architecture.
+As illustrated in Fig.
+2, we first establish a Mamba-based image restoration baseline using an asymmetric encoder-decoder design within a U-shaped architecture.
 The restoration process begins with a $3\times 3$ convolutional layer, which extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded image.
@@
 Specifically, we propose a Degradation-Aware Prompt State Space Model, which guides the image restoration process by compressing historical information and supplementing global degradation information.
-3.4. Pretrained Degradation Extractor
-Figure 3. (a) Degradation Modeling Process of Real-World Images, (b) The Architecture of Degradation Extractor.
-Figure 4. The Architecture of (a) Degradation-aware Prompt Selective Scan (DPSS) and (b) High-frequency Enhancement Block (HEB), with Three Inputs $F_{1}$, $F_{2}$, and Degradation Embedding $D$.
-To better extract degradation information from images, we revisit the degradation modeling process of real-world images, which spans from image capture and imaging to transmission, as illustrated in Figure  3 (a).
+III-C Degradation Extractor
+To better extract degradation information from images, we revisit the degradation modeling process of real-world images, which spans from image capture and imaging to transmission, as illustrated in Fig.
+3 (a).
 This process typically involves four common types of degradations: (1) poor illumination conditions, (2) relative motion between the target objects and the camera, (3) artifacts introduced by transmission media in the optical path, and (4) compression during storage and transportation.
-We extend these four types of degradations to a more general framework and propose a unified degradation imaging model.
-The specific formulation is defined as follows:
-(6) $I(x)=Q\big{(}(\alpha(x)J(x)+\beta(x)\gamma(x))\otimes K\big{)}.$
+We unify these scenarios into a generalized degradation formation model:
+$I(x)=Q\big((\alpha(x)J(x)+\beta(x)\gamma(x))\otimes K\big).$ (5)
 where $J(x)$ represents the clean image, $\alpha(x)$ models the luminance adjustment, such as the effects of insufficient lighting in low-light or adverse weather conditions (e.g., rain or fog), and $\gamma(x)$ accounts for artifacts caused by occlusions from aerosols, raindrops, or snow particles in the transmission medium.
 $\beta(x)$ modulates $\gamma(x)$ based on varying spatial properties, $K$ represents the blur introduced by relative motion between the target and the camera, and $Q$ denotes the compression applied during storage and transmission.
-Based on the aforementioned degradation modeling process, we design a universal degradation extractor, as illustrated in Figure 3 (b).
-Specifically, we introduce central difference convolution(CDC)(Wang et al., 2023a; Peng et al., 2024a, b) to extract edge and texture degradation cues related to illumination and blur.
-In parallel, we employ three groups of CDC with different scales to capture artifacts of varying sizes, such as visual artifacts and snowflakes.
+Figure 4: The Architecture of (a) Degradation-aware Prompt Selective Scan (DPSS) and (b) High-frequency Enhancement Block (HEB), with Three Inputs $F_{1}$, $F_{2}$, and Degradation Embedding $\boldsymbol{E}_{d}$.
+Based on the aforementioned degradation modeling process, we design a universal degradation extractor, as illustrated in Fig.
+3 (b).
+Specifically, we introduce central difference convolution (CDC)[36, 37, 38] to extract edge and texture degradation cues related to illumination and blur.
+In addition, we propose a multi-scale degradation perception module, which employs four parallel CDC operators of different sizes to capture degradation features at various scales.
+By concatenating the outputs, the module retains complete degradation representations across all frequency components.
+Furthermore, we employ three groups of CDC with different scales to capture artifacts of varying sizes, such as visual artifacts and snowflakes.
 Additionally, we integrate a traditional convolution branch to learn brightness-related representations.
-Furthermore, to train the degradation extractor with fine-grained degradation encoding capability, we propose a pretraining method for the degradation extractor based on degraded image reconstruction.
-This method leverages a hybrid dataset consisting of paired data with diverse types of degradation.
-The core of this process is training the degradation extractor $\mathcal{E}$ to extract fine-grained degradation information from the degraded image $\boldsymbol{I}_{D}$ and distill it into a compact degradation embedding $D$ of size $1\times 512$, formulated as $D=\mathcal{E}(\boldsymbol{I}_{D})$.
-Subsequently, the reconstruction unit $\mathcal{R}$ utilizes $D$ and the clean image $\boldsymbol{O}$ to reconstruct the degraded image $\boldsymbol{I}_{D}$, formulated as $\boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{O},D)$.
+To achieve a balance between extraction performance and computational cost, we introduce a re-parameterization technique for efficient integration of degradation representations, as illustrated in Fig.
+3 (c).
+These two implementations correspond to different variants of our method, which are presented in Table I as DPMambaIR-L and DPMambaIR, respectively.
+Furthermore, to endow the degradation extractor with fine-grained degradation encoding capability, we propose a pretraining strategy based on degraded image reconstruction.
+This strategy leverages a hybrid dataset composed of paired samples with diverse degradation types.
+The key objective is to train the degradation extractor $\mathcal{E}$ to capture detailed degradation information from a degraded image $\boldsymbol{I}_{D}$ and compress it into a compact embedding $\boldsymbol{E}_{d}$, formulated as:
+$\boldsymbol{E}_{d}=\mathcal{E}(\boldsymbol{I}_{D}).$ (6)
+Subsequently, the reconstruction unit $\mathcal{R}$ takes $\boldsymbol{E}_{d}$ together with the clean content embedding $\boldsymbol{E}_{c}$ to reconstruct the degraded image $\boldsymbol{I}_{D}$.
+In particular, we adopt Restormer as the content extractor to derive $\boldsymbol{E}_{c}$ from the clean reference image $\boldsymbol{O}$, expressed as:
+$\boldsymbol{E}_{c}=Restormer(\boldsymbol{O}),\ \ \boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{E}_{c},\boldsymbol{E}_{d}).$ (7)
+This pretraining paradigm encourages the degradation extractor to encode both the type and severity of degradation into a continuous embedding space, which then serves as a compact and informative task descriptor for downstream restoration.
 We further introduce a dual-component loss function $L$, which combines the $L_{1}$ norm and the Learned Perceptual Image Patch Similarity (LPIPS) metric.
-The combined loss function is expressed as $L(\boldsymbol{I}_{D},\boldsymbol{\hat{I}}_{D})=L_{1}+L_{LPIPS}$, striking a careful balance between reconstruction accuracy and perceptual quality.
-3.5. Degradation-aware Prompt State Space Model
-To address these limitation mentioned in 3.2, we propose a novel Degradation-aware Prompt State Space Model to fully unlock the potential of SSM in All-in-One image restoration.
+The combined loss function is expressed as $L(\boldsymbol{I}_{D},\boldsymbol{\hat{I}}_{D})=L_{1}+L_{LPIPS}$, striking a balance between reconstruction accuracy and perceptual quality.
+III-D Degradation-aware Prompt State Space Model
+To address the limitations of SSM-based image restoration—specifically the insufficient use of global context and the lack of effective degradation-aware mechanisms, we propose a novel Degradation-aware Prompt State Space Model (DP-SSM) to fully unlock the potential of SSM in All-in-One image restoration.
 Specifically, we propose a novel Degradation-aware Prompt Selective Scan, which incorporates degradation information into the State Space Model.
@@
 In addition, we modulate the parameter matrices in the state-space equations to establish a degradation-aware mechanism.
-This mechanism enables SSM to adapt to the imbalances caused by the inherent differences among sub-tasks in All-in-One image restoration.
-To further mitigate the issue of high-frequency information loss, which arises from balancing the differences among sub-tasks during training, we design a simple yet effective High-frequency Enhancement Block (HEB).
-This component ensures the preservation of high-frequency details and significantly enhances the overall performance of the model.
+This mechanism enables SSM to adapt to multi-task learning and handle task competition.
+To better recover local high-frequency details, we design a lightweight High-frequency Enhancement Block (HEB).
+This component ensures the preservation of high-frequency details and enhances the overall model performance.
 Degradation-aware Prompt Selective Scan.
-We revisit the traditional SSM, whose parameter settings are shown in Equation (3).
+We revisit the traditional SSM, whose parameter settings are shown in Eq.(2).
 Matrices $A$, $B$, $C$, and $D$ control the SSM output.
@@
 In traditional SSM, these matrices and $\Delta$ are derived from:
-(7) $A=init(),D=init(),$
+$A=init(),D=init(),$ (8)
 $B=Linear(x),C=Linear(x),\Delta=Linear(x).$
 where $init()$ is an initialization method that does not depend on the input $x$.
-This way may lead to two key issues: (a) Traditional State Space Model struggles to effectively leverage global contextual information.
-(b) The lack of a proper degradation-aware mechanism makes it challenging to balance the differences among tasks in multi-task learning.
-To address these issues, we first utilize a pre-trained degradation extractor $\mathcal{E}$ to extract the degradation embedding $D$ from the input image $I_{D}$.
-Subsequently, $D$ is used to modulate $\Delta$, the input matrix $B$, and the output matrix $C$ of the SSM.
-This introduces a degradation-aware mechanism in three aspects: historical information compression, input information processing, and global context awareness, as illustrated in Figure 4 (a).
-The specific method is formulated as follows:
-(8) $D$ $=\mathcal{E}(I_{D}),$
-$\alpha_{\Delta}$ $=\mathcal{M}_{\Delta}(D),\,\alpha_{B}=\mathcal{M}_{B}(D),\,\alpha_{C}=\mathcal{M}_{C}(D),$
+However, such parameterization does not incorporate degradation information, which motivates us to introduce a degradation-aware prompt mechanism to guide the state-space modeling.
+To integrate degradation awareness into the SSM, we first utilize a degradation extractor $\mathcal{E}$ to extract the degradation embedding $\boldsymbol{E}_{d}$ from the input image $I_{D}$.
+This embedding is then employed to modulate $\Delta$, the input matrix $B$, and the output matrix $C$, thereby enhancing the model’s ability to compress historical information, process input signals, and capture global context, as illustrated in Fig.
+4 (a).
+The specific formulation is given as follows:
+$\boldsymbol{E}_{d}$ $=\mathcal{E}(I_{D}),$ (9)
+$\alpha_{\Delta}$ $=\mathcal{M}_{\Delta}(\boldsymbol{E}_{d}),\,\alpha_{B}=\mathcal{M}_{B}(\boldsymbol{E}_{d}),\,\alpha_{C}=\mathcal{M}_{C}(\boldsymbol{E}_{d}),$
 $\Delta_{dp}$ $=\alpha_{\Delta}\cdot\Delta,\,B_{dp}=\alpha_{B}\cdot B,\,C_{dp}=\alpha_{C}\cdot C.$
-Here, $\mathcal{M}_{\Delta}$, $\mathcal{M}_{B}$, and $\mathcal{M}_{C}$ are combinations of linear layers used to map the degradation embedding $\mathbf{e}$ into feature spaces, forming modulation vectors $\alpha_{\Delta}$, $\alpha_{B}$, and $\alpha_{C}$ for $\Delta$, $B$, and $C$, respectively.
+Here, $\mathcal{M}_{\Delta}$, $\mathcal{M}_{B}$, and $\mathcal{M}_{C}$ are combinations of linear layers used to map the degradation embedding $\boldsymbol{E}_{d}$ into feature spaces, forming modulation vectors $\alpha_{\Delta}$, $\alpha_{B}$, and $\alpha_{C}$ for $\Delta$, $B$, and $C$, respectively.
 $\Delta_{dp}$, $B_{dp}$, and $C_{dp}$ are the newly modulated SSM parameters.
-Finally, the state-space function formulation of the Degradation-aware State Space Model can be expressed as follows:
-(9) $\overline{A}_{dp}$ $=\exp(\Delta_{dp}A),$
+Finally, the formulation of the Degradation-aware State Space Model can be expressed as follows:
+$\overline{A}_{dp}$ $=\exp(\Delta_{dp}A),$ (10)
 $\overline{B}_{dp}$ $=(\Delta_{dp}A)^{-1}(\exp(A)-I)\cdot\Delta_{dp}B_{dp},$
@@
 The output matrix $C$ controls how historical information affects the output.
-In traditional SSM, the $i$-th pixel can only observe the $i-1$ preceding pixels, leading to a limited global receptive field.
-While existing methods address this with parallel multi-directional scanning, this increases computational costs and offers limited performance gains in low-level vision tasks (Guo et al., 2024c).
+In traditional SSM, the $i$-th pixel can only observe the i1i-1 preceding pixels, leading to a limited global receptive field.
+While existing methods address this with parallel multi-directional scanning, this increases computational costs and offers limited performance gains in low-level vision tasks [32].
 Instead, we retain single-directional scanning and enhance global degradation cues via $C$, achieving efficient global awareness with lower computational overhead.
 High-frequency Enhancement Block.
-Figure 5.
-We compare frequency-domain differences between clear and restored images under three configurations.
+Figure 5: We compare frequency-domain differences between clear and restored images under three configurations.
 The All-in-One setting shows a notable high-frequency gap compared to the Degradation-specific setting, which is significantly reduced by adding the High-frequency Enhancement Block (HEB), as shown by the blue line.
-In All-in-One image restoration tasks, models tend to prioritize low-frequency restoration over high-frequency details, such as edges and textures, leading to their loss, as shown in Figure 5.
-To address this, we propose an efficient high-frequency enhancement block with minimal computational overhead.
-For input features $F$, we first extract channel-wise low-frequency features using global average pooling, then subtract them from $F$ to obtain high-frequency features, as shown in Figure 4(b).
+All-in-One image restoration involves recovering image components distributed across different frequency bands.
+While our proposed Degradation-aware Prompt State Space Model enables dynamic adaptation to degradation characteristics and partially alleviates inter-task conflicts, it still exhibits a bias toward low-frequency restoration due to the inherent global modeling nature of Mamba.
+As shown in Fig.
+5, All-in-One models tend to underperform in reconstructing high-frequency details such as edges and textures, compared to degradation-specific models.
+To address this limitation, we introduce an extremely lightweight High-Frequency Enhancement Block (HEB) as a complementary module.
+For input features $F$, we first extract channel-wise low-frequency features using global average pooling, then subtract them from $F$ to obtain high-frequency components, as illustrated in Fig.
+4(b).
 Finally, the high-frequency features are added back to enhance $F$, formulated as:
-(10) ${F}_{enhanced}={F}+\alpha\cdot\big{(}{F}-\mathcal{G}({F})\big{)}.$
-Here, $\alpha$ represents a learnable parameter that controls the weight of high-frequency information during the enhancement process.
-$\mathcal{G}$ represents the global average pooling operation.
-${F}_{enhanced}$ denotes the features after high-frequency enhancement.
-3.6. Loss Function
+${F}_{enhanced}={F}+\alpha\cdot\big({F}-\mathcal{G}({F})\big).$ (11)
+Here, $\alpha$ represents a learnable scalar that controls the contribution of high-frequency enhancement, and $\mathcal{G}$ denotes the global average pooling operation.
+${F}_{enhanced}$ is the enhanced feature map.
+III-E Loss Function
 We employ the commonly used L1 loss $\mathcal{L}_{1}$ and L2 loss $\mathcal{L}_{2}$ in low-level vision tasks.
-Additionally, following previous works(Zamir et al., 2022; Cui et al., 2023), we incorporate a frequency-domain loss $\mathcal{L}_{fft}$ for training.
+Additionally, following previous works [39, 40], we incorporate a frequency-domain loss $\mathcal{L}_{fft}$ for training.
 The total loss is defined as:
-(11) $\mathcal{L}_{total}=\lambda_{1}\cdot\mathcal{L}_{1}(O,\hat{O})+\lambda_{2}\cdot\mathcal{L}_{2}(O,\hat{O})+\lambda_{3}\cdot\mathcal{L}_{fft}(O,\hat{O}).$
+[…]
 where $O$ and $\hat{O}$ denote the model output and the ground truth, respectively.
 The parameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are balancing factors, which we set to 1, 0.5 and 0.001, respectively.
-4. Experiment And Analysis
+Figure 6: Examples of Seven Types of Image Degradations in the Dataset.
+TABLE I: Quantitative Comparison on Seven Datasets for All-in-One Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
+Method Lowlight Snow Rain Haze Jpeg Noise Blur Average PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM MPRNet 19.71 0.659 30.94 0.938 26.27 0.912 25.63 0.947 30.51 0.938 28.03 0.897 26.07 0.869 26.74 0.880 MIRNet 19.81 0.668 31.04 0.941 25.77 0.907 22.26 0.882 30.54 0.939 28.14 0.900 26.32 0.873 26.27 0.873 NAFNet 19.49 0.650 30.33 0.935 26.04 0.906 25.63 0.946 30.20 0.936 27.80 0.895 26.13 0.870 26.52 0.877 Restormer 19.99 0.670 31.67 0.944 26.91 0.919 26.18 0.953 30.56 0.939 28.17 0.890 26.62 0.880 27.16 0.886 MambaIR 19.81 0.664 30.43 0.935 25.52 0.901 25.90 0.944 30.38 0.937 27.91 0.895 26.07 0.869 26.57 0.878 PromptIR 19.90 0.669 31.52 0.943 26.80 0.918 26.13 0.945 30.57 0.939 28.18 0.901 26.36 0.874 27.07 0.884 IDR 19.68 0.663 31.16 0.941 26.39 0.914 26.11 0.948 30.51 0.939 28.10 0.899 26.29 0.873 26.89 0.883 NDR-Restore 19.79 0.635 30.00 0.902 26.00 0.813 27.00 0.943 30.19 0.872 27.78 0.786 25.77 0.791 26.64 0.820 OneRestore 19.77 0.657 29.95 0.931 25.37 0.900 27.74 0.962 30.10 0.934 27.65 0.890 25.90 0.866 26.64 0.877 MoCEIR 19.92 0.674 31.22 0.943 26.56 0.920 28.06 0.963 30.51 0.939 28.13 0.900 26.34 0.874 27.25 0.888 AdaIR 19.99 0.672 31.94 0.923 27.01 0.922 26.61 0.952 30.53 0.939 28.19 0.900 26.61 0.879 27.26 0.887 DPMambaIR 20.04 0.680 32.36 0.949 27.07 0.926 28.13 0.963 30.57 0.939 28.25 0.902 27.43 0.894 27.69 0.893 DPMambaIR-L 20.06 0.677 32.31 0.950 27.11 0.925 28.39 0.963 30.55 0.939 28.21 0.902 27.45 0.895 27.73 0.893
+TABLE II: Quantitative Comparison on Deraining, Low-light Image Enhancement(LLIE), Deblurring and Dehazing for Degradation-Specific Image Restoration Task with State-of-the-Art Methods.
+The Best and Second-Best Results Are in Bold and Underline.
+Deraining PSNR SSIM JORDER 26.25 0.835 PReNet 29.46 0.899 MPRNet 30.41 0.891 MAXIM 30.81 0.903 Restormer 31.46 0.904 OneRestore 29.36 0.944 MoCEIR 31.23 0.960 AdaIR 31.46 0.961 DPMambaIR 32.30 0.967
+LLIE PSNR SSIM EnlightenGAN 17.61 0.653 MIRNet 24.14 0.830 URetinex-Net 19.84 0.824 MAXIM 23.43 0.863 IAT 23.38 0.809 OneRestore 22.97 0.835 MoCEIR 23.64 0.818 AdaIR 23.47 0.831 DPMambaIR 24.20 0.852
+Deblurring PSNR SSIM DeepDeBlur 29.08 0.913 DeBlurGAN 28.70 0.858 DeBlurGANV2 29.55 0.934 MT-RNN 31.15 0.945 IR-SDE 30.70 0.901 OneRestore 28.76 0.915 MoCEIR 30.05 0.933 AdaIR 30.57 0.939 DPMambaIR 31.42 0.948
+Dehazing PSNR SSIM GCANet 26.59 0.935 GridDehazeNet 25.86 0.944 DeHazeFormer 30.29 0.964 MAXIM 29.12 0.932 DA-CLIP 30.16 0.936 OneRestore 30.16 0.974 MoCEIR 30.62 0.975 AdaIR 30.87 0.975 DPMambaIR 30.89 0.977
+IV Experiment and Analysis
 We evaluate our method on both All-in-one image restoration tasks and Degradation-specific image restoration tasks.
@@
 For Degradation-specific image restoration, we train separately on datasets corresponding to each type of degradation.
-4.1. Experimental Settings
+IV-A Experimental Settings
 Training Details.
-Building on previous works(Zamir et al., 2022; Potlapalli et al., 2023; Guo et al., 2024c), we develop a Mamba-based baseline for image restoration.
-Our method is trained using the AdamW(Kingma and Ba, 2014) optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $1e^{-4}$ for 300,000 iterations.
+Building on previous works[39, 22, 32], we develop a Mamba-based baseline for image restoration.
+Our method is trained using the AdamW[41] optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $1e^{-4}$ for 300,000 iterations.
 The initial learning rate of $3e^{-4}$ is gradually reduced to $1e^{-6}$ through cosine annealing.
-Following(Zamir et al., 2022), we employ a progressive training strategy, cropping image patches of size $192\times 192$.
+Following[39], we employ a progressive training strategy, cropping image patches of size $192\times 192$.
 Horizontal and vertical flips are applied for data augmentation.
@@
 Datasets.
-For All-in-One image restoration, we collect a large-scale mixed dataset that includes seven classic and popular degradation types: fog, rain, snow, low light, blur, JPEG compression artifacts, and noise.
-Additionally, we evaluate performance on Degradation-specific tasks, using the LOL dataset(Wei et al., 2018) for low-light image enhancement, the GoPro dataset(Nah et al., 2017a) for image deblurring, the Rain100H dataset(Yang et al., 2017) for image deraining, and the RESIDE6K dataset(Qin et al., 2020) for image dehazing.
-Table 1. Quantitative Comparison on Seven Datasets for All-in-One Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
-Method Lowlight Snow Rain Fog Jpeg Noise Blur Average PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM MPRNet 19.71 0.659 30.94 0.938 26.27 0.912 25.63 0.947 30.51 0.938 28.03 0.897 26.07 0.869 26.74 0.880 MIRNet 19.81 0.668 31.04 0.941 25.77 0.907 22.26 0.882 30.54 0.939 28.14 0.900 26.32 0.873 26.27 0.873 NAFNet 19.49 0.650 30.33 0.935 26.04 0.906 25.63 0.946 30.20 0.936 27.80 0.895 26.13 0.870 26.52 0.877 Restormer 19.99 0.670 31.67 0.944 26.91 0.919 26.18 0.953 30.56 0.939 28.17 0.890 26.62 0.880 27.16 0.886 PromptIR 19.90 0.669 31.52 0.943 26.80 0.918 26.13 0.945 30.57 0.939 28.18 0.901 26.36 0.874 27.07 0.884 IDR 19.68 0.663 31.16 0.941 26.39 0.914 26.11 0.948 30.51 0.939 28.10 0.899 26.29 0.873 26.89 0.883 OneRestore 19.77 0.657 29.95 0.931 25.37 0.900 27.74 0.962 30.10 0.934 27.65 0.890 25.90 0.866 26.64 0.877 AdaIR 19.99 0.672 31.94 0.923 27.01 0.922 26.61 0.952 30.53 0.939 28.19 0.900 26.61 0.879 27.26 0.887 DPMambaIR 20.04 0.680 32.36 0.949 27.07 0.926 28.13 0.963 30.57 0.939 28.25 0.902 27.43 0.894 27.69 0.893
-Table 2.
-Quantitative Comparison on Deraining, Low-light Image Enhancement(LLIE), Deblurring and Dehazing for Degradation-Specific Image Restoration Task with State-of-the-Art Methods.
-The Best and Second-Best Results Are in Bold and Underline.
-Deraining PSNR SSIM JORDER 26.25 0.835 PReNet 29.46 0.899 MPRNet 30.41 0.891 MAXIM 30.81 0.903 Restormer 31.46 0.904 DPMambaIR 32.30 0.967
-LLIE PSNR SSIM EnlightenGAN 17.61 0.653 MIRNet 24.14 0.830 URetinex-Net 19.84 0.824 MAXIM 23.43 0.863 IAT 23.38 0.809 DPMambaIR 24.20 0.852
-Deblurring PSNR SSIM DeepDeBlur 29.08 0.913 DeBlurGAN 28.70 0.858 DeBlurGANV2 29.55 0.934 MT-RNN 31.15 0.945 IR-SDE 30.70 0.901 DPMambaIR 31.42 0.948
-Dehazing PSNR SSIM GCANet 26.59 0.935 GridDehazeNet 25.86 0.944 DeHazeFormer 30.29 0.964 MAXIM 29.12 0.932 DA-CLIP 30.16 0.936 DPMambaIR 30.89 0.977
-Figure 6. Qualitative comparison on Deblurring, Deraining and Dehazing for All-in-One Image Restoration Task.
+For All-in-One image restoration, we collect a large-scale mixed dataset covering seven common degradation types: haze, rain, snow, low-light, blur, noise, and JPEG compression, as shown in Fig.
+6.
+Detailed information regarding these datasets is summarized in Table III.
+Additionally, degradation-specific evaluations are conducted using LOL [42] for low-light enhancement, GoPro [43] for deblurring, Rain100H [44] for deraining, and RESIDE6K [45] for dehazing.
+TABLE III: Datasets for Image Restoration Tasks
+Degradation Dataset Train Pairs Test Pairs
+Haze RESIDE6K [45] 6000 1000
+Rain Rain13K [46] 13711 -
+Rain100H [44] - 100
+Snow Snow100K [47] 5000 1000
+Low Light LOL [42], LSRW [48] 6085 65
+Blur GoPro [43] 2103 1111
+Noise WaterlooED [49] 4744 -
+BSD400 [50] - 400
+JPEG WaterlooED [49] 4744 -
+BSD400 [50] - 400
+Figure 7: Qualitative comparison on Deblurring, Deraining and Dehazing for All-in-One Image Restoration Task.
 Evaluation Metrics.
-Following previous works(Zamir et al., 2022; Guo et al., 2024c), we use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively compare different methods on the RGB channels.
+Following previous works[39, 32], we use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively compare different methods on the RGB channels.
 Comparisons with State-of-the-art Methods.
-For All-in-One image restoration, we compare our method against eight approaches: MIRNet (Zamir et al., 2020), NAFNet (Chen et al., 2022), MPRNet (Zamir et al., 2021), Restormer (Zamir et al., 2022), PromptIR (Potlapalli et al., 2023), IDR (Zhang et al., 2023a), OneRestore (Guo et al., 2024a), and AdaIR (Cui et al., 2025).
-Among these, MIRNet, NAFNet, MPRNet, and Restormer are general-purpose image restoration models, while the latter four methods are specifically designed for All-in-One image restoration.
-Specifically, PromptIR employs visual prompts to guide restoration, OneRestore leverages a pre-trained degradation category encoder with cross-attention mechanisms, and AdaIR utilizes frequency domain information for adaptive restoration.
-For Degradation-specific image restoration tasks, we compare our method with different state-of-the-art approaches tailored to each domain.
-The details are as follows: (1) For deraining, we compare against JORDER (Yang et al., 2019), PReNet (Ren et al., 2019), MPRNet (Zamir et al., 2021), MAXIM (Tu et al., 2022), and Restormer (Zamir et al., 2022).
-(2) For low-light image enhancement, we compare against EnlightenGAN (Jiang et al., 2021), MIRNet (Zamir et al., 2020), Uretinex-Net (Wu et al., 2022), MAXIM (Tu et al., 2022), and IAT (Cui et al., 2022).
-(3) For deblurring, we compare against DeepDeblur (Nah et al., 2017b), DeblurGAN (Kupyn et al., 2017), DeblurGANv2 (Kupyn et al., 2019), MT-RNN (Park et al., 2020), and IR-SDE (Luo et al., 2023b).
-(4) For dehazing, we compare against GCANet (Chen et al., 2018), GridDehazeNet (Liu et al., 2019), DeHazeFormer (Song et al., 2023), MAXIM (Tu et al., 2022), and DA-CLIP (Luo et al., 2023a).
-4.2. Quantitative Results
+For All-in-One image restoration, we compare against eleven methods: MIRNet [51], NAFNet [52], MPRNet [53], Restormer [39], MambaIR [32], PromptIR [22], IDR [23], NDR-Restore [54], OneRestore [9], MoCEIR [55], and AdaIR [24].
+The first four are general-purpose restoration models, while the latter four are designed specifically for All-in-One restoration: PromptIR uses visual prompts, OneRestore employs a pre-trained degradation encoder with cross-attention, and AdaIR leverages frequency-domain information for adaptive restoration.
+For degradation-specific restoration, we also include representative single-task methods such as JORDER [56] (deraining), MAXIM [57] (deblurring), EnlightenGAN [58] (low-light enhancement), DeepDeblur [59], AdaIR [24], DA-CLIP [25], among others.
+IV-B Quantitative Results
 All-in-one Image Restoration.
-We compare our method against five state-of-the-art image restoration approaches, including both CNN-based and transformer-based methods.
-As shown in Table 1, for the All-in-One image restoration task, our method achieves the best PSNR and SSIM on the aforementioned mixed datasets compared to existing methods.
-Specifically, our method achieves a PSNR of 27.69dB and an SSIM of 0.893, outperforming AdaIR by 0.43dB and 0.011, respectively.
+We conducted two versions of our method for comparison: the normal version (DPMambaIR) and a Large version (DPMambaIR-L).
+The normal version adopts the degradation extractor illustrated in Fig.
+3(c), while the Large version uses the architecture shown in Fig.
+3(b).
+For subsequent evaluations, we primarily adopt the normal version unless otherwise specified.
+We compare our method against eleven state-of-the-art image restoration approaches.
+As shown in Table I, for the All-in-One image restoration task, our method achieves the best PSNR and SSIM on the aforementioned mixed datasets, outperforming existing methods.
+Specifically, our method achieves a PSNR of 27.69dB and an SSIM of 0.893, surpassing AdaIR by 0.43dB and 0.011, respectively.
+Additionally, DPMambaIR-L achieves a PSNR of 27.73dB and an SSIM of 0.893, slightly outperforming the normal version.
 We achieve the best performance across all seven sub-tasks.
@@
 Degradation-Specific Image Restoration.
-We evaluate our method on four Degradation-specific tasks, consistently achieving superior performance across multiple datasets compared to state-of-the-art CNN-based, transformer-based, and diffusion-based approaches (Table 2).
-On Rain100H (Yang et al., 2017), our method achieves a PSNR of 32.30dB and SSIM of 0.967, surpassing Restormer by 0.84dB and 0.063, respectively.
-On the LOL dataset (Wei et al., 2018), it achieves 24.20dB PSNR and 0.852 SSIM, outperforming MIRNet by 0.06dB and 0.022.
-For deblurring on GoPro (Nah et al., 2017a), it attains 31.42dB PSNR and 0.948 SSIM, exceeding MT-RNN by 0.27dB and 0.003.
-On RESIDE6K (Qin et al., 2020), it achieves 30.89dB PSNR and 0.973 SSIM, outperforming DehazeFormer by 0.60dB and 0.009.
+We evaluate our method on four Degradation-specific tasks, consistently achieving superior performance across multiple datasets compared to state-of-the-art CNN-based, transformer-based, and diffusion-based approaches, as shown in Table II.
+On Rain100H [44], our method achieves a PSNR of 32.30dB and SSIM of 0.967, surpassing AdaIR by 0.84dB and 0.006, respectively.
+On the LOL dataset [42], it achieves 24.20dB PSNR and 0.852 SSIM, outperforming MIRNet by 0.06dB and 0.022.
+For deblurring on GoPro [43], it attains 31.42dB PSNR and 0.948 SSIM, exceeding MT-RNN by 0.27dB and 0.003.
+On RESIDE6K [45], it achieves 30.89dB PSNR and 0.973 SSIM, outperforming AdaIR by 0.02dB and 0.002.
 These results highlight the versatility and effectiveness of our method in addressing diverse degradation restoration tasks.
-4.3. Qualitative Results
-Figure 6 presents the visual results of deblurring, deraining, and dehazing under the All-in-One Image Restoration task setting.
+IV-C Qualitative Results
+Fig.
+7 presents the visual results of deblurring, deraining, and dehazing under the All-in-One Image Restoration task setting.
 Our method demonstrates superior visual performance compared to other approaches.
@@
 Additional visual comparisons under both the All-in-One and Degradation-specific settings are provided in the Appendix.
-4.4. Ablation Study
-Table 3 presents the results of our ablation study on DPMambaIR for the All-in-One image restoration task, illustrating the impact of removing the degradation prompt, different modulation modes, and the High-frequency Enhancement Block (HEB) from the model.
-For the baseline model without the Degradation-Aware Prompt Selective Scan and HEB, the PSNR and SSIM are only 26.96 dB and 0.884, respectively.
-Introducing degradation-awareness via prompts applied to Matrices $A$ and $B$ improves PSNR and SSIM by 0.64 dB and 0.008, respectively, by enabling degradation-aware mechanisms in historical information compression and input feature processing.
-Adding Prompt $C$, which supplements global degradation information, further increases PSNR by 0.07 dB.
-The inclusion of HEB significantly boosts performance, preserving high-frequency details.
-The complete model achieves 27.69 dB PSNR and 0.893 SSIM, demonstrating the importance of all components for optimal performance.
-Table 4 presents the results for degradation extraction and utilization methods.
-For degradation extraction, we compare three approaches: explicit category labels, a pre-trained degradation extractor from OneRestore (Guo et al., 2024a), and our fine-grained degradation representation via image reconstruction.
-The OneRestore extractor, designed for classification tasks, performs worse than explicit category labels.
-In contrast, our method achieves a 0.11 dB PSNR improvement over non-blind restoration with category labels by leveraging fine-grained degradation representations.
-For degradation utilization, we compare concatenation and attention, with our method outperforming both, demonstrating its effectiveness and superiority in extracting and utilizing degradation information.
-For the ablation study on global information supplementation, we compare our method with Mamba’s dual-sequence scanning scheme.
-As shown in Table 5, our approach not only achieves superior performance compared to the dual-sequence scanning method but also demonstrates a significant reduction in both the Parameters and FLOPs.
-Table 3. Ablation Studies on our proposed core Modules.
-[…]
-Table 4. Ablation Study on Different Degradation Extraction and Utilization Methods.
-Comparison Projects Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Ours 27.69 0.893
-Table 5. Ablation Study on Different Methods for Global Information Supplementation.
-Method PSNR SSIM Params(M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt Matrix C 27.69 0.893 29.35 153.3
-5. Conclusion
-In this paper, we proposed DPMambaIR, a novel All-in-One framework for image restoration that balances high restoration quality and computational efficiency across diverse degradation types.
-DPMambaIR introduces a Degradation-Aware Prompt State Space Model (DP-SSM), which dynamically integrates fine-grained degradation features into the state-space modeling process.
-This design enables adaptive handling of complex degradation scenarios while maintaining global degradation awareness.
-Furthermore, the High-Frequency Enhancement Block (HEB) reduces high-frequency detail loss caused by task competition, ensuring superior edge and texture restoration.
+IV-D Ablation Study
+We conduct ablation studies to evaluate our key designs: the degradation-aware prompts and High-frequency Enhancement Block (HEB) (Table IV), degradation extraction and utilization methods (Table V), and the global information supplementation strategy (Table VI).
+Table IV shows the effectiveness of our core modules.
+The High-frequency Enhancement Block (HEB) alone improves the baseline by 0.29 dB PSNR by enhancing local details.
+Degradation-aware prompts ($P_{\Delta}$, $P_{B}$) bring a more significant gain of 0.64 dB, showing the benefit of adaptive modulation.
+Adding $P_{C}$ for global cues offers further slight improvement.
+The full model demonstrates the synergy between prompt-based modulation and frequency enhancement, achieving the best performance.
+TABLE IV: Ablation Studies on Our Proposed Core Modules.
+Method $P_{\Delta}$ $P_{B}$ $P_{C}$ HEB PSNR SSIM Baseline 26.96 0.884 (a) 27.25 0.887 (b) 27.60 0.892 (c) 27.67 0.893 (d) 27.69 0.893
+Figure 8: t-SNE visualization of degradation embeddings.
+Table V presents the results for degradation extraction and utilization methods.
+For degradation extraction, we compare three approaches: (1) “Classes”, which provides manually specified degradation types and can be regarded as a 100% accurate non-blind setting; (2) a pre-trained degradation classifier from OneRestore [9]; and (3) our fine-grained degradation representation via image reconstruction.
+The OneRestore extractor, optimized for classification, is insensitive to degradation strength or location, leading to poorer results.
+Our method surpasses the non-blind ”Classes” baseline by 0.11 dB PSNR.
+Fig.
+8 visualizes the t-SNE embeddings of the extracted degradation features.
+Significant gaps exist between different degradation types, indicating clear separability.
+Notably, although JPEG and noise share the same content during synthesis, the t-SNE visualization reveals that their separability is primarily driven by degradation information rather than image content.
+For degradation utilization, we compare concatenation and attention-based fusion with our prompt modulation design.
+Our prompt-based modulation outperforms simple concatenation or attention-based fusion, showing better adaptability to local degradations.
+TABLE V: Ablation Study on Different Degradation Extraction and Utilization Methods.
+Comparison Projects Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Ours 27.69 0.893 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Ours 27.69 0.893
+Table VI presents the ablation study on global information supplementation.
+We compare our Prompt-C with Mamba’s bi-directional scanning scheme.
+Compared to Mamba’s bi-directional scan, our Prompt-C not only achieves slightly better performance but is also more efficient, reducing GFLOPs by nearly 37% (Table VI).
+This highlights its effectiveness as a lightweight global information supplementation method.
+TABLE VI: Ablation Study on Different Methods for Global Information Supplementation.
+Method PSNR SSIM Params(M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt Matrix C (Ours) 27.69 0.893 29.35 153.3
+V Conclusion
+We propose DPMambaIR, a novel All-in-One framework for image restoration capable of handling diverse degradation types.
+The core of our framework is a Degradation-Aware Prompt State Space Model (DP-SSM) that leverages a fine-grained degradation extractor.
+This design enables the dynamic integration of degradation features into the state-space modeling process, allowing for adaptive handling of complex degradation scenarios while maintaining global degradation awareness.
+Furthermore, a lightweight High-frequency Enhancement Block (HEB) was introduced to complement the main framework by enhancing high-frequency details restoration with negligible computational overhead.
 This study underscores the importance of fine-grained degradation modeling and dynamic feature modulation in advancing All-in-One image restoration frameworks.

### v3

@@
 Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs.
-Recent advances in deep learning have significantly improved performance in Degradation-specific restoration tasks, such as image denoising [1, 2], deblurring [3], deraining[4], dehazing[5, 6] and low-light enhancement[7].
+Recent advances in deep learning have significantly improved performance in degradation-specific restoration tasks, such as image denoising [64, 19], deblurring [66], deraining[38], dehazing[43, 47], and low-light enhancement[30].
 However, traditional methods typically rely on degradation-specific models explicitly tailored to individual degradation types.
@@
 Existing methods can be broadly categorized into Mixture-of-Experts (MoE)-based and Prompt-based approaches.
-MoE-based methods, such as MoFME [8], dynamically route tasks to specialized experts using an uncertainty-aware router, enabling effective learning of diverse expert functions.
-However, these methods suffer from high computational costs, increased inference complexity as the number of experts grows, and lack explicit degradation modeling, limiting adaptability to unseen or complex degradations.
-Prompt-based methods, like OneRestore [9], guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms for flexible and adaptive strategies.
-While effective, these methods rely on coarse prompts that fail to capture nuanced degradation characteristics and treat prompts as static inputs, limiting their ability to adapt to varying degradation intensities or specific requirements, leading to suboptimal results.
-While these methods have demonstrated success in learning multiple tasks for All-in-One image restoration, two key challenges remain unresolved.
-First, existing models lack a degradation extractor capable of capturing fine-grained degradation features and often rely on coarse degradation labels to guide the restoration process.
-For example, images with varying degrees of haze are typically assigned a single haze label, failing to capture the nuanced variations in degradation levels.
-This simplification leads to suboptimal restoration performance, as the model cannot adapt to the diverse characteristics of different degradations.
-Second, current methods struggle to balance the competing requirements of varying restoration tasks.
-For instance, training low-light enhancement, which focuses on restoring low-frequency contrast information, alongside deblurring, which prioritizes the recovery of fine local details.
-This inherent task competition makes it challenging to optimize a single network for all restoration tasks simultaneously and remains a fundamental challenge for All-in-One image restoration frameworks.
-To address these challenges, we propose DPMambaIR, a novel All-in-One image restoration framework built upon a regression-based degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM).
-Unlike prior approaches that use static or class-based prompts, our method employs a degradation extractor to generate prompt vectors that encode both semantic categories and degradation distribution characteristics.
-These prompts are then used to dynamically modulate the state transition, input, and output matrices of a state space model, enabling fine-grained degradation modeling and adaptive coordination across frequency bands.
-This mechanism allows our model to better manage task competition and enhance restoration quality across diverse degradations.
-Moreover, while State Space Model such as Mamba exhibit strong global modeling capabilities, they are inherently biased toward low-frequency restoration [10].
-To better recover local high-frequency details, we introduce a lightweight High-Frequency Enhancement Block (HEB) as a supplementary component, which delivers performance gains while incurring only negligible computational overhead.
-Extensive experiments on a mixed dataset containing seven degradation types demonstrate the superiority of DPMambaIR.
+MoE-based methods, such as MoFME [63], dynamically route tasks to specialized experts using an uncertainty-aware router.
+However, these methods suffer from high computational costs and lack explicit degradation modeling, limiting adaptability to unseen degradations.
+Alternatively, Prompt-based methods like OneRestore [11] guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms.
+However, existing prompt-based strategies [11, 39] suffer from fundamental limitations in how degradation information is represented and exploited.
+A key shortcoming lies in their reliance on static priors, where degradations are encoded as discrete class labels or fixed semantic tokens.
+Such categorical representations are inherently inadequate for modeling the continuous nature of real-world degradations and their spatial variability, for example, differentiating between light haze and dense, spatially accumulating fog.
+Furthermore, most of these methods follow an additive fusion paradigm in which degradation cues are treated merely as auxiliary information appended to visual features.
+This mechanism is intrinsically limited, as the underlying network architecture remains governed by fixed convolutional kernels or static attention patterns.
+Consequently, the model is compelled to address heterogeneous restoration demands, such as suppressing high-frequency noise while simultaneously enhancing global contrast, within a rigid geometric structure.
+This conflict ultimately degrades performance across diverse and complex degradation scenarios.
+To address these challenges, we propose DPMambaIR, a novel All-in-One framework that shifts the design philosophy from conventional feature fusion to explicit dynamic parameter modulation.
+Our approach is built upon a fine-grained degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM).
+Unlike methods that rely on discrete degradation labels, the proposed extractor employs a reconstruction objective to regress continuous degradation embeddings, enabling it to encode nuanced information regarding both degradation type and severity.
+These embeddings are utilized to directly modulate the core State Space Model parameters $(\Delta,B,C)$.
+Modulating the discretization step $\Delta$ effectively alters the integration step size of the underlying ordinary differential equation, allowing the model to realize distinct continuous dynamics conditioned on degradation characteristics.
+Specifically, the network learns to assign smaller $\Delta$ values to induce a high-inertia state favorable for smoothing high-frequency noise, whereas larger values are generated to enter a high-gain regime that amplifies weak signals in low-light conditions.
+Concurrently, the input projection matrix $B$ is modulated to control how strongly external observations drive the latent state transitions, while the output matrix $C$ is adjusted to inject global degradation priors.
+Furthermore, we observe that jointly optimizing heterogeneous degradations often induces an implicit bias toward low-frequency structures, making the consistent recovery of fine details challenging.
+To alleviate this issue, we incorporate a lightweight High-frequency Enhancement Block (HEB) that complements the proposed dynamic modulation and facilitates the restoration of local textures.
+Extensive experiments on a mixed dataset containing seven degradation types demonstrate the efficacy of DPMambaIR.
 As shown in Fig.
 1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore.
-Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, showcasing its potential as an effective and robust All-in-One image restoration solution.
+Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, demonstrating its potential as an effective and robust All-in-One image restoration solution.
 The contributions of this paper are summarized as follows:
   • $\bullet$
-    We propose a novel degradation extractor capable of capturing fine-grained degradation features from complex degraded images.
-These features provide critical priors to improve the performance of All-in-One image restoration.
+    We propose a degradation extractor capable of capturing fine-grained, continuous degradation features from complex degraded images via a regression-based reconstruction objective.
   • $\bullet$
-    We design a Degradation-aware Prompt State Space Model (DP-SSM), which enhances the long-range modeling capability of Mamba by integrating degradation priors at multiple structural levels.
+    We design a Degradation-aware Prompt State Space Model (DP-SSM) that introduces a dynamic parameter modulation mechanism.
+By dynamically modulating the state-space evolution parameters based on degradation priors, it enhances the model’s physical adaptability to diverse degradation dynamics.
   • $\bullet$
     Extensive experiments on a mixed dataset containing seven types of degradation demonstrate that DPMambaIR achieves state-of-the-art performance on PSNR and SSIM metrics, outperforming existing methods.
-The results validate the effectiveness and superiority of the proposed approach.
+The results validate the effectiveness and robustness of the proposed approach.
 II Related Work
 II-A All-in-One Image Restoration
-Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs.
-Traditional methods, such as Dark Channel Prior [11] and Color Line Prior [12], rely on handcrafted priors to model degradation but struggle with complex scenarios due to predefined assumptions.
-The rise of deep learning [13] has significantly advanced tasks like such as image super-resolution[14, 15, 16], image denoising [1, 2], deblurring [3], deraining[4], dehazing[5], desnowing[17] and low-light enhancement[7].
-These approaches, leveraging convolutional neural networks (CNN)-based [18], transformer-based [19], or diffusion-based [20] architectures, enable powerful feature learning for image restoration.
-However, most existing works focus on addressing a single type of degradation, which limits their applicability in real-world scenarios where images often suffer from diverse and unknown types of degradation.
-Recent advancements in All-in-One image restoration frameworks focus on enhancing flexibility and efficiency by leveraging various multi-task learning techniques such as Mixture-of-Experts (MoE) and text-based or vision-based prompts to handle diverse restoration tasks.
-For example, AirNet[21] employs contrastive learning to extract degradation representations for guiding restoration.
-PromptIR[22] integrates end-to-end prompt generation and interaction to adaptively handle different degradation types, while OneRestore[9] aligns text-based degradation categories with visual features as prompts.
-IDR[23] uses meta-learning to decompose degradations, enabling generalization across diverse scenarios.
-MoFME[8], on the other hand, employs an uncertainty-aware router to dynamically assign Degradation-specific features to expert modules, allowing it to specialize in multiple tasks.
-Similarly, AdaIR[24] achieves adaptive restoration by decoupling degradations in the frequency domain.
-Additionally, methods like DA-CLIP[25] leverage the pre-trained vision-language model (VLM) for prompt extraction, but the significant computational overhead introduced by VLM limits their practical efficiency.
-Despite their effectiveness, most prompt-based methods rely on coarse-grained class labels or static degradation tokens, which limits their ability to model continuous degradation distributions or adapt to image-specific degradation characteristics.
-In contrast, our method extracts fine-grained, regression-level degradation embeddings that encode both semantic and statistical degradation features.
-These embeddings are used to dynamically modulate the state-space parameters of the restoration backbone, enabling degradation-aware processing throughout the network.
-This facilitates better coordination across restoration tasks and reduces competition between frequency components.
+Image restoration remains a foundational challenge in computer vision, striving to reconstruct high-fidelity content from corrupted inputs.
+While traditional approaches, such as Dark Channel Prior [13] and Color Line Prior [7], relied on handcrafted heuristics that often struggle in complex scenarios, the advent of deep learning [21] has revolutionized the field.
+Significant strides have been made in specific domains, including super-resolution [24, 36, 37], denoising [64, 19], deblurring [66, 61, 62], deraining [38, 59], dehazing [43, 17], desnowing [10, 60], and low-light enhancement [30].
+Leveraging CNN-based [22], Transformer-based [45], or diffusion-based [14] architectures, these methods excel at feature representation.
+However, their reliance on degradation-specific designs fundamentally limits their utility in real-world environments where degradations are diverse, compounded, and often unknown.
+Recent research has consequently pivoted towards All-in-One frameworks that emphasize flexibility and unified processing.
+These methods generally employ multi-task learning strategies, such as Mixture-of-Experts (MoE) or prompt-guided adaptation.
+For instance, AirNet [23] utilizes contrastive learning to derive degradation representations, while MoFME [63] implements an uncertainty-aware router to dynamically dispatch features to specialized experts.
+Although effective, MoE architectures often incur high computational overhead and lack explicit physical modeling of the degradation.
+In parallel, prompt-based methods such as PromptIR [39], OneRestore [11], and DA-CLIP [27] integrate degradation cues, ranging from learned visual prompts to text embeddings, via cross-attention mechanisms.
+IDR [58] further explores meta-learning for decomposition, and AdaIR [4] adopts frequency domain decoupling for adaptive restoration.
+Despite these advancements, fundamental limitations persist in the representation and exploitation of degradation information.
+Most prompt-based strategies [39, 11] rely on coarse-grained priors or discrete class labels, which fail to capture the continuous nature and spatial variability inherent in real-world images.
+Furthermore, they typically adopt an additive fusion paradigm where degradation cues are treated merely as auxiliary inputs.
+This mechanism is intrinsically limited because the underlying network parameters remain fixed, governed by static weights or attention patterns that lack the flexibility to adapt to heterogeneous restoration demands.
+In contrast, our DPMambaIR framework employs a regression-based extractor to derive fine-grained, continuous degradation embeddings.
+We depart from simple feature concatenation and utilize these embeddings to dynamically re-parameterize the State Space Model.
+This shift from additive fusion to explicit dynamic parameter modulation allows the system to adapt its evolution dynamics to the specific physical properties of the degradation.
 Figure 2: The overall framework of DPMambaIR consists of a shallow feature extractor and multiple DPSS blocks, with an upsampler implemented by pixel shuffle and a downsampler by pixel unshuffle.
@@
 II-B State Space Model based Image Restoration
-In Image Restoration, Convolutional Neural Networks (CNN) [13] and Transformer [19] have been widely adopted due to their strong feature extraction and representation learning capabilities.
-CNN excels at capturing local patterns, making them suitable for tasks requiring high spatial resolution and fine details.
-However, their limited receptive field often hinders their ability to model long-range dependencies, which are critical for handling global degradations.
-On the other hand, Transformer leverages self-attention mechanisms to capture global context effectively.
-Despite their success, Transformers suffer from quadratic complexity relative to the input size, making them computationally and memory-intensive.
-To address these limitations, researchers have turned to the State Space Model (SSM)[26], which provides an efficient framework for modeling long-range dependencies with linear computational complexity.
-Formally, SSM maps an input sequence $x(t)$ to an output sequence $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$, represented by a linear ordinary differential equation:
+Convolutional Neural Networks (CNNs) [21] and Transformers [45] have long dominated image restoration.
+While CNNs excel at capturing local high-frequency details, their limited receptive fields hinder the modeling of long-range dependencies essential for correcting global degradations like haze or uneven lighting.
+Transformers address this with global self-attention but suffer from quadratic computational complexity relative to image resolution.
+To reconcile long-range modeling with efficiency, researchers have revisited the State Space Model (SSM) [18].
+SSMs map a 1D input sequence $x(t)$ to an output $y(t)$ via a latent state $h(t)\in\mathbb{R}^{N}$, governed by the linear ordinary differential equation (ODE):
 $h^{\prime}(t)$ $=Ah(t)+Bx(t),$ (1)
 $y(t)$ $=Ch(t)+Dx(t).$
-where $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}$.
-After discretization with step size $\Delta$, the dynamic equation can be expressed as:
+where $A,B,C,D$ are learnable parameters. For digital implementation, this continuous system is discretized with a step size $\Delta$, yielding the recurrence:
 $h_{i}$ $=\overline{A}h_{i-1}+\overline{B}x_{i},$ (2)
 $y_{i}$ $=Ch_{i}+Dx_{i}.$
-where $\overline{A}=\exp(\Delta A)$ and $\overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B$.
-The introduction of Vision Mamba[27] marked the first effort to incorporate SSM into vision tasks.
-By proposing the Vision State Space Module (VSSM), it extended SSM to visual applications and demonstrated superior performance over Vision Transformer (ViT)[28], while maintaining lower complexity.
-This success inspired extensive research into SSM across diverse tasks such as object detection[29], semantic segmentation[30], and image classification[31].
-In the context of Image Restoration, MambaIR[32] pioneered the integration of SSM, leveraging its linear complexity and long-range modeling capability to achieve significant improvements in both computational efficiency and perceptual quality.
-This was followed by several variants targeting specific restoration tasks[33, 34, 35], collectively showcasing the strong potential of SSM in this domain.
-Existing approaches typically unfold 2D images into 1D sequences and apply sequential causal modeling.
-However, this design still faces two limitations: (a) Limited Global Context Utilization: The $i$-th pixel can only access the previous i1i-1 pixels in the sequence, restricting the effective use of global context.
-Parallel multi-directional scanning has been explored to enlarge the receptive field, but it increases computational cost while providing limited gains in low-level vision tasks[32].
-(b) Lack of Effective Degradation-Aware Mechanisms: Current SSM-based methods primarily target single-task restoration.
-Their transition and output mappings are fixed or solely input-driven, without leveraging external degradation priors.
-As a result, they cannot dynamically adapt to diverse degradation types or levels, leading to suboptimal performance in All-in-One image restoration task.
-These limitations motivate the introduction of a degradation-aware modulation mechanism into the SSM pipeline.
-In this work, we propose a Degradation-aware Prompt State Space Model (DP-SSM), where degradation embeddings dynamically modulate the state-space transition and output matrices.
-This design enables a fine-grained, task-aware mechanism that balances global modeling with degradation-specific guidance, enhancing the flexibility and effectiveness of SSM for complex restoration scenarios.
+Here, the discretized matrices are defined as $\overline{A}=\exp(\Delta A)$ and $\overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B$.
+This formulation allows for linear-complexity inference while maintaining a global receptive field.
+The Vision Mamba [65] first adapted SSMs for computer vision, demonstrating that the Vision State Space Module could outperform Vision Transformers [6] with significantly lower computational costs.
+This success has sparked adoption across tasks including object detection [50], segmentation [49], and classification [42].
+In image restoration, MambaIR [9] pioneered the use of SSMs to achieve a balance between efficiency and perceptual quality, inspiring subsequent variants [8, 35, 5].
+However, applying standard SSMs to All-in-One restoration faces two primary limitations.
+First, Limited Global Context Utilization: The causal nature of standard scanning means the $i$-th pixel only accesses information from preceding pixels.
+While multi-directional scanning can alleviate this, it increases computation without always yielding proportional gains in low-level tasks [9].
+Second, and more critically, Lack of Degradation-Awareness: Existing SSM-based restoration methods rely on fixed transition dynamics ($A,B,C$) or purely input-driven modulation.
+They lack an explicit mechanism to adapt the system’s evolution rules based on external degradation priors (e.g., adjusting the integration step $\Delta$ for different noise levels).
+Consequently, their ability to generalize across the diverse degradation types encountered in All-in-One settings remains constrained.
+Motivated by these gaps, we propose the Degradation-aware Prompt State Space Model (DP-SSM).
+Unlike prior works, our method utilizes fine-grained degradation embeddings to dynamically modulate the discretization step $\Delta$, along with the state-space transition and output matrices.
+This design establishes a dynamics-aware mechanism that harmonizes global context modeling with degradation-specific physical guidance, significantly enhancing robustness in complex restoration scenarios.
 III Method
@@
 The degradation process $Deg$ can be mathematically represented as:
-$\boldsymbol{I}_{D}=Deg(\boldsymbol{I}_{C})+\boldsymbol{\eta}.$ (4)
-Here, $\boldsymbol{I}_{C}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $Deg$ represents the degradation operator (e.g., blur kernel, brightness, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise.
+$\boldsymbol{I}_{D}=Deg(\boldsymbol{O})+\boldsymbol{\eta}.$ (4)
+Here, $\boldsymbol{O}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $Deg$ represents the degradation operator (e.g., blur kernel, brightness, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise.
 The main challenge arises from the fact that $Deg$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
@@
 As illustrated in Fig.
-2, we first establish a Mamba-based image restoration baseline using an asymmetric encoder-decoder design within a U-shaped architecture.
-The restoration process begins with a $3\times 3$ convolutional layer, which extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded image.
-These features then pass through three encoding stages, each followed by downsampling.
-During each encoding stage, the resolution of the feature map is halved while the number of channels is doubled.
-Subsequently, the features pass through three decoding stages, where each stage involves upsampling and concatenating the output of the decoder with the corresponding output of the encoder to form the input for the next decoding stage.
-Finally, after a refinement stage, the clean image is restored through a shortcut connection that combines the input degraded image with the output of the decoder.
-To address the challenges posed by unknown degradation types and levels in All-in-One image restoration tasks, we propose a Degradation Extractor pre-trained using degraded image reconstruction.
-This model compresses degradation type and severity information into a compact embedding, enabling the model to transition from blind to non-blind image restoration through finer-grained degradation representation.
-To further optimize the utilization of the extracted degradation embedding, we revisit the selective state-space modeling process.
-Specifically, we propose a Degradation-Aware Prompt State Space Model, which guides the image restoration process by compressing historical information and supplementing global degradation information.
+2, our method is built upon a Mamba-based U-shaped architecture with an asymmetric encoder-decoder design.
+Initially, a $3\times 3$ convolution extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded input.
+The encoder comprises three stages, each utilizing downsampling to progressively halve the feature resolution while doubling the channel dimension.
+Correspondingly, the decoder employs upsampling and skip connections to fuse multi-scale features from the encoding path.
+Finally, the clean image is reconstructed via a refinement stage, utilizing a global residual connection to superimpose the learned residual onto the degraded input.
+To address the challenges posed by unknown degradation types and severities in All-in-One image restoration, we propose a Degradation Extractor pre-trained via a self-supervised reconstruction objective.
+This module encodes heterogeneous degradation patterns into a compact embedding, effectively bridging the gap between blind and non-blind restoration by providing explicit degradation representations.
+To fully exploit these representations, we revisit the selective state-space modeling paradigm.
+Specifically, we propose a Degradation-Aware Prompt State Space Model, which dynamically modulates the restoration trajectory by injecting global degradation priors into the state transition process.
 III-C Degradation Extractor
-To better extract degradation information from images, we revisit the degradation modeling process of real-world images, which spans from image capture and imaging to transmission, as illustrated in Fig.
+To effectively extract degradation priors, we formulate a generalized degradation model that encapsulates the physical corruption process from image capture to transmission, as shown in Fig.
 3 (a).
-This process typically involves four common types of degradations: (1) poor illumination conditions, (2) relative motion between the target objects and the camera, (3) artifacts introduced by transmission media in the optical path, and (4) compression during storage and transportation.
-We unify these scenarios into a generalized degradation formation model:
+This process typically involves four degradation categories: (1) illumination degradation, (2) motion-induced blur, (3) transmission medium artifacts, and (4) compression artifacts.
+We unify these factors into a composite formation model:
 $I(x)=Q\big((\alpha(x)J(x)+\beta(x)\gamma(x))\otimes K\big).$ (5)
-where $J(x)$ represents the clean image, $\alpha(x)$ models the luminance adjustment, such as the effects of insufficient lighting in low-light or adverse weather conditions (e.g., rain or fog), and $\gamma(x)$ accounts for artifacts caused by occlusions from aerosols, raindrops, or snow particles in the transmission medium.
-$\beta(x)$ modulates $\gamma(x)$ based on varying spatial properties, $K$ represents the blur introduced by relative motion between the target and the camera, and $Q$ denotes the compression applied during storage and transmission.
+where $J(x)$ denotes the latent clean image.
+The term $\alpha(x)$ models luminance variations, encompassing low-light conditions or uneven illumination caused by weather.
+$\gamma(x)$ represents additive artifacts introduced by the transmission medium, such as rain streaks, snow, or haze aerosols, while $\beta(x)$ modulates the spatial intensity of these artifacts.
+$K$ denotes the blur kernel resulting from relative motion, and $Q$ represents quantization or compression operations applied during storage.
+Guided by this formulation, we design a degradation extractor capable of capturing heterogeneous corruption patterns, as illustrated in Fig.
+3(b).
+Specifically, we employ Central Difference Convolution (CDC) [46, 33, 34] to identify gradient-based degradation cues, such as edges and textures associated with blur and noise.
+A multi-scale module aggregates features from four parallel CDC branches with varying kernel sizes, enabling the perception of degradations across different frequency bands.
+Additionally, we integrate a standard convolution branch to capture global intensity and brightness distributions.
+To balance extraction accuracy with computational efficiency, we employ a re-parameterization technique during inference, merging the multi-branch structure into a compact representation, as shown in Fig.
+3 (c).
+This yields two model variants: a larger capacity DPMambaIR-L utilizing the full multi-branch structure, and a lightweight DPMambaIR employing the re-parameterized extractor (detailed in Table I).
+To enable the learning of fine-grained degradation embeddings, we propose a reconstruction-based pre-training strategy.
+Unlike classification-based methods that rely on discrete labels, this approach forces the extractor to learn a continuous manifold of degradation types and severities.
+Formally, the extractor $\mathcal{E}$ maps a degraded input $\boldsymbol{I}_{D}$ to a low-dimensional embedding $\boldsymbol{E}_{d}$:
+$\boldsymbol{E}_{d}=\mathcal{E}(\boldsymbol{I}_{D}).$ (6)
+To supervise this process, a reconstruction decoder $\mathcal{R}$ attempts to recover the degraded image $\boldsymbol{I}_{D}$ by combining $\boldsymbol{E}_{d}$ with a content embedding $\boldsymbol{E}_{c}$ derived from the corresponding clean ground truth $\boldsymbol{O}$:
+$\boldsymbol{E}_{c}=\text{Restormer}(\boldsymbol{O}),\ \ \boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{E}_{c},\boldsymbol{E}_{d}).$ (7)
+This disentanglement ensures that $\boldsymbol{E}_{d}$ captures pure degradation information necessary to reconstruct the corruption, independent of the image content.
+The optimization minimizes a hybrid objective combining $L_{1}$ loss and the Learned Perceptual Image Patch Similarity (LPIPS) metric, defined as $\mathcal{L}=\|\boldsymbol{I}_{D}-\boldsymbol{\hat{I}}_{D}\|_{1}+\lambda\mathcal{L}_{\text{LPIPS}}$, to ensure perceptually faithful reconstruction of the degradation patterns.
 Figure 4: The Architecture of (a) Degradation-aware Prompt Selective Scan (DPSS) and (b) High-frequency Enhancement Block (HEB), with Three Inputs $F_{1}$, $F_{2}$, and Degradation Embedding $\boldsymbol{E}_{d}$.
-Based on the aforementioned degradation modeling process, we design a universal degradation extractor, as illustrated in Fig.
-3 (b).
-Specifically, we introduce central difference convolution (CDC)[36, 37, 38] to extract edge and texture degradation cues related to illumination and blur.
-In addition, we propose a multi-scale degradation perception module, which employs four parallel CDC operators of different sizes to capture degradation features at various scales.
-By concatenating the outputs, the module retains complete degradation representations across all frequency components.
-Furthermore, we employ three groups of CDC with different scales to capture artifacts of varying sizes, such as visual artifacts and snowflakes.
-Additionally, we integrate a traditional convolution branch to learn brightness-related representations.
-To achieve a balance between extraction performance and computational cost, we introduce a re-parameterization technique for efficient integration of degradation representations, as illustrated in Fig.
-3 (c).
-These two implementations correspond to different variants of our method, which are presented in Table I as DPMambaIR-L and DPMambaIR, respectively.
-Furthermore, to endow the degradation extractor with fine-grained degradation encoding capability, we propose a pretraining strategy based on degraded image reconstruction.
-This strategy leverages a hybrid dataset composed of paired samples with diverse degradation types.
-The key objective is to train the degradation extractor $\mathcal{E}$ to capture detailed degradation information from a degraded image $\boldsymbol{I}_{D}$ and compress it into a compact embedding $\boldsymbol{E}_{d}$, formulated as:
-$\boldsymbol{E}_{d}=\mathcal{E}(\boldsymbol{I}_{D}).$ (6)
-Subsequently, the reconstruction unit $\mathcal{R}$ takes $\boldsymbol{E}_{d}$ together with the clean content embedding $\boldsymbol{E}_{c}$ to reconstruct the degraded image $\boldsymbol{I}_{D}$.
-In particular, we adopt Restormer as the content extractor to derive $\boldsymbol{E}_{c}$ from the clean reference image $\boldsymbol{O}$, expressed as:
-$\boldsymbol{E}_{c}=Restormer(\boldsymbol{O}),\ \ \boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{E}_{c},\boldsymbol{E}_{d}).$ (7)
-This pretraining paradigm encourages the degradation extractor to encode both the type and severity of degradation into a continuous embedding space, which then serves as a compact and informative task descriptor for downstream restoration.
-We further introduce a dual-component loss function $L$, which combines the $L_{1}$ norm and the Learned Perceptual Image Patch Similarity (LPIPS) metric.
-The combined loss function is expressed as $L(\boldsymbol{I}_{D},\boldsymbol{\hat{I}}_{D})=L_{1}+L_{LPIPS}$, striking a balance between reconstruction accuracy and perceptual quality.
 III-D Degradation-aware Prompt State Space Model
-To address the limitations of SSM-based image restoration—specifically the insufficient use of global context and the lack of effective degradation-aware mechanisms, we propose a novel Degradation-aware Prompt State Space Model (DP-SSM) to fully unlock the potential of SSM in All-in-One image restoration.
+To address the limitations of SSM-based image restoration, specifically the insufficient use of global context and the lack of effective degradation-aware mechanisms, we propose a novel Degradation-aware Prompt State Space Model (DP-SSM) to effectively adapt SSM to All-in-One image restoration.
 Specifically, we propose a novel Degradation-aware Prompt Selective Scan, which incorporates degradation information into the State Space Model.
@@
 where $init()$ is an initialization method that does not depend on the input $x$.
-However, such parameterization does not incorporate degradation information, which motivates us to introduce a degradation-aware prompt mechanism to guide the state-space modeling.
+However, this input-dependent design is agnostic to the degradation type.
+It processes high-frequency noise and low-frequency blur with the same initialization prior, lacking the ability to structurally adapt the system dynamics for specific restoration tasks.
+This motivates the introduction of a degradation-aware prompt mechanism, which guides the state-space modeling by incorporating specific degradation cues.
 To integrate degradation awareness into the SSM, we first utilize a degradation extractor $\mathcal{E}$ to extract the degradation embedding $\boldsymbol{E}_{d}$ from the input image $I_{D}$.
@@
 $\overline{A}_{dp}$ $=\exp(\Delta_{dp}A),$ (10)
-$\overline{B}_{dp}$ $=(\Delta_{dp}A)^{-1}(\exp(A)-I)\cdot\Delta_{dp}B_{dp},$
+$\overline{B}_{dp}$ $=(\Delta_{dp}A)^{-1}(\exp(\Delta_{dp}A)-I)\cdot\Delta_{dp}B_{dp},$
 $h_{i}$ $=\overline{A}_{dp_{i}}h_{i-1}+\overline{B}_{dp_{i}}x_{i},$
 $y_{i}$ $=C_{dp}h_{i}+Dx_{i}.$
-Through this approach, the transformation matrix $A$ and input matrix $B$ gain degradation-aware capabilities, integrating degradation cues of various types and levels during historical information compression and input signal processing, making the model more task-friendly.
-The output matrix $C$ controls how historical information affects the output.
-In traditional SSM, the $i$-th pixel can only observe the i1i-1 preceding pixels, leading to a limited global receptive field.
-While existing methods address this with parallel multi-directional scanning, this increases computational costs and offers limited performance gains in low-level vision tasks [32].
-Instead, we retain single-directional scanning and enhance global degradation cues via $C$, achieving efficient global awareness with lower computational overhead.
-High-frequency Enhancement Block.
+The proposed modulation mechanism fundamentally alters the underlying system dynamics, effectively recalibrating the trade-off between memory inertia and input sensitivity.
+Central to this adaptation is the step size $\Delta_{dp}$, whose learned distribution shown in Fig.
+5 reveals a distinct physical correspondence to the degradation characteristics.
+For high-frequency degradations such as noise and JPEG artifacts, the model autonomously converges to a minimal $\Delta_{dp}$ regime.
+As $\Delta_{dp}$ approaches zero, the discretized transition matrix $\overline{{A}}_{dp}$ approximates the identity matrix while the input gain $\overline{{B}}_{dp}$ diminishes.
+This configuration transforms the SSM into a stable low-pass filter, creating a high-inertia state that effectively suppresses instantaneous pixel variance by prioritizing historical context over the unreliable current observation.
+In contrast, inputs characterized by low-frequency or structural deficiencies, such as blur and low-light conditions, trigger a marked increase in the learned $\Delta_{dp}$.
+This shift transitions the system into a high-gain regime, amplifying the magnitude of $\overline{{B}}_{dp}$ to enhance sensitivity towards weak input signals and facilitate the recovery of attenuated local structures.
+Complementing this dynamic discretization, the concurrent modulation of B{B} acts as a gating mechanism for input observations, while the modulation of C{C} supplements the feature readout with global degradation cues.
+Figure 5: Distribution of the learned parameter $\Delta_{dp}$ across different degradation types in the first 20% of DPSS layers.
+The model adaptively assigns smaller step sizes for noise/jpeg and larger step sizes for blur/low-light, confirming physically interpretable dynamic modulation.
+Through this strategy, the transformation matrix A{A} and input matrix B{B} are endowed with degradation-aware capabilities, allowing them to adaptively regulate memory inertia and input sensitivity according to the specific degradation severity.
+However, the standard SSM remains constrained by its strictly causal nature, where the $i$-th pixel can only access information from the preceding i1i-1 pixels, resulting in a restricted global receptive field.
+Although existing methods attempt to alleviate this via parallel multi-directional scanning, such approaches incur high computational costs and yield limited performance gains in low-level vision tasks [9].
+By retaining the efficient single-directional scanning and instead modulating matrix C{C} to inject global degradation information, our method achieves effective global degradation awareness with significantly lower computational overhead.
 Figure 6: We compare frequency-domain differences between clear and restored images under three configurations.
 The All-in-One setting shows a notable high-frequency gap compared to the Degradation-specific setting, which is significantly reduced by adding the High-frequency Enhancement Block (HEB), as shown by the blue line.
-All-in-One image restoration involves recovering image components distributed across different frequency bands.
-While our proposed Degradation-aware Prompt State Space Model enables dynamic adaptation to degradation characteristics and partially alleviates inter-task conflicts, it still exhibits a bias toward low-frequency restoration due to the inherent global modeling nature of Mamba.
-As shown in Fig.
-5, All-in-One models tend to underperform in reconstructing high-frequency details such as edges and textures, compared to degradation-specific models.
-To address this limitation, we introduce an extremely lightweight High-Frequency Enhancement Block (HEB) as a complementary module.
+High-frequency Enhancement Block.
+All-in-One image restoration necessitates the recovery of image components distributed across diverse frequency bands.
+While our proposed Degradation-aware Prompt State Space Model enables dynamic adaptation to degradation characteristics and partially mitigates inter-task conflicts, the joint optimization of heterogeneous degradations inevitably induces an implicit bias toward low-frequency structures.
+As illustrated in Fig.
+6, this spectral bias results in suboptimal performance regarding high-frequency details, such as edges and textures, compared to degradation-specific models.
+To address this limitation, we introduce a lightweight High-Frequency Enhancement Block (HEB) as a complementary module to explicitly facilitate the restoration of local textures.
 For input features $F$, we first extract channel-wise low-frequency features using global average pooling, then subtract them from $F$ to obtain high-frequency components, as illustrated in Fig.
@@
 […]
-where $O$ and $\hat{O}$ denote the model output and the ground truth, respectively.
+where $\hat{O}$ and $O$ denote the model output and the ground truth, respectively.
 The parameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are balancing factors, which we set to 1, 0.5 and 0.001, respectively.
 Figure 7: Examples of Seven Types of Image Degradations in the Dataset.
-TABLE I: Quantitative Comparison on Seven Datasets for All-in-One Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
-Method Lowlight Snow Rain Haze Jpeg Noise Blur Average PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM MPRNet 19.71 0.659 30.94 0.938 26.27 0.912 25.63 0.947 30.51 0.938 28.03 0.897 26.07 0.869 26.74 0.880 MIRNet 19.81 0.668 31.04 0.941 25.77 0.907 22.26 0.882 30.54 0.939 28.14 0.900 26.32 0.873 26.27 0.873 NAFNet 19.49 0.650 30.33 0.935 26.04 0.906 25.63 0.946 30.20 0.936 27.80 0.895 26.13 0.870 26.52 0.877 Restormer 19.99 0.670 31.67 0.944 26.91 0.919 26.18 0.953 30.56 0.939 28.17 0.890 26.62 0.880 27.16 0.886 MambaIR 19.81 0.664 30.43 0.935 25.52 0.901 25.90 0.944 30.38 0.937 27.91 0.895 26.07 0.869 26.57 0.878 PromptIR 19.90 0.669 31.52 0.943 26.80 0.918 26.13 0.945 30.57 0.939 28.18 0.901 26.36 0.874 27.07 0.884 IDR 19.68 0.663 31.16 0.941 26.39 0.914 26.11 0.948 30.51 0.939 28.10 0.899 26.29 0.873 26.89 0.883 NDR-Restore 19.79 0.635 30.00 0.902 26.00 0.813 27.00 0.943 30.19 0.872 27.78 0.786 25.77 0.791 26.64 0.820 OneRestore 19.77 0.657 29.95 0.931 25.37 0.900 27.74 0.962 30.10 0.934 27.65 0.890 25.90 0.866 26.64 0.877 MoCEIR 19.92 0.674 31.22 0.943 26.56 0.920 28.06 0.963 30.51 0.939 28.13 0.900 26.34 0.874 27.25 0.888 AdaIR 19.99 0.672 31.94 0.923 27.01 0.922 26.61 0.952 30.53 0.939 28.19 0.900 26.61 0.879 27.26 0.887 DPMambaIR 20.04 0.680 32.36 0.949 27.07 0.926 28.13 0.963 30.57 0.939 28.25 0.902 27.43 0.894 27.69 0.893 DPMambaIR-L 20.06 0.677 32.31 0.950 27.11 0.925 28.39 0.963 30.55 0.939 28.21 0.902 27.45 0.895 27.73 0.893
+TABLE I: Quantitative Comparison on Seven Datasets for All-in-One Image Restoration.
+The best and second-best results are shown in bold and underline, respectively.
+(Format: PSNR $\uparrow$  / SSIM $\uparrow$  / LPIPS $\downarrow$).
+Method Lowlight Snow Rain Haze Jpeg Noise Blur Average MPRNet 19.71 / .659 / .122 30.94 / .938 / .029 26.27 / .912 / .068 25.63 / .947 / .023 30.51 / .938 / .024 28.03 / .897 / .057 26.07 / .869 / .077 26.74 / .880 / .057 MIRNet 19.81 / .668 / .123 31.04 / .941 / .028 25.77 / .907 / .073 22.26 / .882 / .058 30.54 / .939 / .024 28.14 / .900 / .057 26.32 / .873 / .070 26.27 / .873 / .062 NAFNet 19.49 / .650 / .117 30.33 / .935 / .030 26.04 / .906 / .068 25.63 / .946 / .023 30.20 / .936 / .025 27.80 / .895 / .051 26.13 / .870 / .075 26.52 / .877 / .056 Restormer 19.99 / .670 / .113 31.67 / .944 / .025 26.91 / .919 / .061 26.18 / .953 / .020 30.56 / .939 / .024 28.17 / .890 / .056 26.62 / .880 / .063 27.16 / .886 / .052 MambaIR 19.81 / .664 / .115 30.43 / .935 / .032 25.52 / .901 / .078 25.90 / .944 / .024 30.38 / .937 / .025 27.91 / .895 / .058 26.07 / .869 / .077 26.57 / .878 / .058 PromptIR 19.90 / .669 / .116 31.52 / .943 / .024 26.80 / .918 / .061 26.13 / .945 / .022 30.57 / .939 / .024 28.18 / .901 / .054 26.36 / .874 / .072 27.07 / .884 / .053 IDR 19.68 / .663 / .115 31.16 / .941 / .027 26.39 / .914 / .065 26.11 / .948 / .023 30.51 / .939 / .024 28.10 / .899 / .054 26.29 / .873 / .073 26.89 / .883 / .054 NDR-Restore 19.79 / .635 / .206 30.00 / .902 / .052 26.00 / .813 / .158 27.00 / .943 / .025 30.19 / .872 / .056 27.78 / .786 / .133 25.77 / .791 / .193 26.64 / .820 / .117 OneRestore 19.77 / .657 / .127 29.95 / .931 / .034 25.37 / .900 / .080 27.74 / .962 / .016 30.10 / .934 / .026 27.65 / .890 / .060 25.90 / .866 / .082 26.64 / .877 / .061 MoCEIR 19.92 / .674 / .109 31.22 / .943 / .025 26.56 / .920 / .051 28.06 / .963 / .014 30.51 / .939 / .023 28.13 / .900 / .053 26.34 / .874 / .069 27.25 / .888 / .049 AdaIR 19.99 / .672 / .102 31.94 / .923 / .023 27.01 / .922 / .056 26.61 / .952 / .019 30.53 / .939 / .024 28.19 / .900 / .055 26.61 / .879 / .065 27.26 / .887 / .049 DPMambaIR 20.04 / .680 / .109 32.36 / .949 / .022 27.07 / .926 / .055 28.13 / .963 / .016 30.57 / .939 / .024 28.25 / .902 / .051 27.43 / .894 / .048 27.69 / .893 / .047 DPMambaIR-L 20.06 / .677 / .108 32.31 / .950 / .022 27.11 / .925 / .056 28.39 / .963 / .016 30.55 / .939 / .024 28.21 / .902 / .051 27.45 / .895 / .045 27.73 / .893 / .046
 TABLE II: Quantitative Comparison on Deraining, Low-light Image Enhancement(LLIE), Deblurring and Dehazing for Degradation-Specific Image Restoration Task with State-of-the-Art Methods.
@@
 Training Details.
-Building on previous works[39, 22, 32], we develop a Mamba-based baseline for image restoration.
-Our method is trained using the AdamW[41] optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $1e^{-4}$ for 300,000 iterations.
-The initial learning rate of $3e^{-4}$ is gradually reduced to $1e^{-6}$ through cosine annealing.
+To ensure effective degradation-aware modeling, we adopt a two-stage training strategy.
+In the first stage, we pre-train the Degradation Extractor using the same mixed-degradation dataset employed for the All-in-One task.
+The training pipeline consists of the Degradation Extractor, a Content Extractor, and a Degradation Reconstructor.
+Specifically, we utilize the pre-trained Restormer from DegAE [25] as the Content Extractor with its parameters frozen, while optimizing only the Degradation Extractor and Reconstructor.
+This pre-training phase is conducted for 100,000 iterations with a batch size of 1, using input patches cropped to $256\times 256$.
+We employ the AdamW [20] optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $1e^{-4}$) with an initial learning rate of $1e^{-4}$, which is gradually reduced to $1e^{-6}$ through cosine annealing.
+In the second stage, the pre-trained Degradation Extractor is integrated into the DPMambaIR framework with its parameters frozen to provide stable guidance.
+Building on previous works [57, 39, 9], we train the main restoration network using the same AdamW configuration for 300,000 iterations.
+The initial learning rate is set to $3e^{-4}$ and decayed to $1e^{-6}$ via cosine annealing.
 Following [57], we employ a progressive training strategy, cropping image patches of size $192\times 192$.
@@
 For All-in-One image restoration, we collect a large-scale mixed dataset covering seven common degradation types: haze, rain, snow, low-light, blur, noise, and JPEG compression, as shown in Fig.
-6.
+7.
 Detailed information regarding these datasets is summarized in Table III.
@@
 Evaluation Metrics.
-Following previous works[39, 32], we use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively compare different methods on the RGB channels.
+Following previous works [57, 9], we employ Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) to quantitatively evaluate restoration quality on RGB channels.
+Additionally, we utilize the Learned Perceptual Image Patch Similarity (LPIPS) to assess the perceptual consistency of the restored images.
 Comparisons with State-of-the-art Methods.
@@
 All-in-one Image Restoration.
-We conducted two versions of our method for comparison: the normal version (DPMambaIR) and a Large version (DPMambaIR-L).
+We evaluated two versions of our method for comparison: the normal version (DPMambaIR) and a large version (DPMambaIR-L).
 The normal version adopts the degradation extractor illustrated in Fig.
@@
 Additional visual comparisons under both the All-in-One and Degradation-specific settings are provided in the Appendix.
-IV-D Ablation Study
-We conduct ablation studies to evaluate our key designs: the degradation-aware prompts and High-frequency Enhancement Block (HEB) (Table IV), degradation extraction and utilization methods (Table V), and the global information supplementation strategy (Table VI).
-Table IV shows the effectiveness of our core modules.
-The High-frequency Enhancement Block (HEB) alone improves the baseline by 0.29 dB PSNR by enhancing local details.
-Degradation-aware prompts ($P_{\Delta}$, $P_{B}$) bring a more significant gain of 0.64 dB, showing the benefit of adaptive modulation.
-Adding $P_{C}$ for global cues offers further slight improvement.
-The full model demonstrates the synergy between prompt-based modulation and frequency enhancement, achieving the best performance.
-TABLE IV: Ablation Studies on Our Proposed Core Modules.
-Method $P_{\Delta}$ $P_{B}$ $P_{C}$ HEB PSNR SSIM Baseline 26.96 0.884 (a) 27.25 0.887 (b) 27.60 0.892 (c) 27.67 0.893 (d) 27.69 0.893
+IV-D Model Complexity and Efficiency Analysis
+To accurately evaluate the practical applicability of DPMambaIR, we conducted a comprehensive efficiency comparison against state-of-the-art All-in-One methods.
+We measured the number of parameters, peak memory consumption, GFLOPs, and average inference time.
+All evaluations were performed on a single NVIDIA RTX A100 GPU with an input tensor size of $1\times 3\times 256\times 256$.
+The statistics for DPMambaIR include the overhead of the Degradation Extractor.
+As summarized in Table IV, DPMambaIR achieves the best performance while maintaining comparable computational overhead.
+Compared to the second-best method, AdaIR, our model delivers a significant performance gain of 0.43 dB with similar parameters and GFLOPs.
+TABLE IV: Comparison of complexity and efficiency against SOTA methods. Metrics: Parameters (M), Peak Memory (MB), GFLOPs, and Inference Time (ms) on $256\times 256$ inputs.
+Method Params (M) Mem.
+(MB) FLOPs (G) Inf Time (ms) PSNR (dB) Restormer 26.13 676.24 154.88 84.283 27.16 IDR 36.28 1433.24 270.76 81.977 26.89 PromptIR 35.59 720.10 172.71 90.964 27.07 MoCEIR 25.35 404.48 97.74 93.232 27.25 AdaIR 28.78 686.40 161.76 102.903 27.26 DPMambaIR 29.35 718.76 153.33 90.869 27.69
+IV-E Ablation Study
+We conduct ablation studies to evaluate our key designs: the degradation-aware prompts and High-frequency Enhancement Block (HEB) (Table V), degradation extraction and utilization methods (Table VI), degradation embedding dimension (Table VII) and the global information supplementation strategy (Table VIII).
+Table V demonstrates the effectiveness of our proposed core modules.
+First, the High-frequency Enhancement Block (HEB) alone improves the baseline by 0.29 dB (Model (a)), validating its efficacy in enhancing local details.
+Second, to verify the importance of the degradation-aware mechanism, we conduct a detailed breakdown of the prompt components.
+Notably, employing $P_{\Delta}$ alone (Model (c)) achieves 27.58 dB, which slightly outperforms the combination of $P_{B}$ and $P_{C}$ (Model (b), 27.56 dB).
+This observation underscores that dynamically modulating the discretization step $\Delta$, which fundamentally alters the state transition dynamics, is the most critical factor for adapting to diverse degradation types.
+Furthermore, combining $P_{\Delta}$ with $P_{B}$ or $P_{C}$ yields consistent improvements, and integrating all prompt modules (Model (f)) pushes the PSNR to 27.67 dB.
+Finally, the full model (Model (g)) demonstrates the synergy between prompt-based modulation and frequency enhancement, achieving the best performance of 27.69 dB.
 Figure 9: t-SNE visualization of degradation embeddings.
-Table V presents the results for degradation extraction and utilization methods.
+Table VI presents the results for degradation extraction and utilization methods.
 For degradation extraction, we compare three approaches: (1) “Classes”, which provides manually specified degradation types and can be regarded as a 100% accurate non-blind setting; (2) a pre-trained degradation classifier from OneRestore [11]; and (3) our fine-grained degradation representation via image reconstruction.
@@
 Our prompt-based modulation outperforms simple concatenation or attention-based fusion, showing better adaptability to local degradations.
-TABLE V: Ablation Study on Different Degradation Extraction and Utilization Methods.
-Comparison Projects Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Ours 27.69 0.893 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Ours 27.69 0.893
-Table VI presents the ablation study on global information supplementation.
+For the training strategy, compared with jointly fine-tuning the degradation extractor together with the backbone network, freezing the parameters of the degradation extractor provides a more stable supply of degradation information.
+This design choice leads to more consistent guidance during restoration and results in a clear performance gain of 0.26 dB in PSNR, demonstrating the advantage of decoupling degradation representation learning from the restoration optimization process.
+TABLE V: Extended ablation studies on the proposed core modules.
+We investigate the individual and combined contributions of the prompt components ($P_{\Delta},P_{B},P_{C}$) and the High-frequency Enhancement Block (HEB).
+Method $P_{\Delta}$ $P_{B}$ $P_{C}$ HEB PSNR SSIM Baseline 26.96 0.884 (a) 27.25 0.887 (b) 27.56 0.891 (c) 27.58 0.892 (d) 27.60 0.892 (e) 27.61 0.892 (f) 27.67 0.893 (g) 27.69 0.893
+TABLE VI: Ablation Study on Different Degradation Extraction and Utilization Methods.
+Module Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Ours 27.69 0.893 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Modulation (Ours) 27.69 0.893 Training Strategy Fine-tuning 27.43 0.888 Frozen (Ours) 27.69 0.893
+TABLE VII: Ablation study on the dimension of degradation embedding ($D_{emb}$).
+We evaluate model complexity (Params, GFLOPs), auxiliary task performance (Rec.
+PSNR, Cls.
+Acc.), and final All-in-One restoration quality (AIO PSNR).
+The default configuration is gray.
+$D_{emb}$ Params FLOPs Rec.
+PSNR Cls.
+Acc.
+AIO.
+PSNR (M) (G) (dB) (%) (dB) 256 28.71 153.33 21.51 72.0 27.55 384 29.03 153.33 24.53 84.3 27.63 512 29.35 153.33 25.67 92.0 27.69 1024 30.62 153.33 25.78 92.6 27.69
+TABLE VIII: Ablation Study on Different Methods for Global Information Supplementation.
+Method PSNR SSIM Params (M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt-C (Ours) 27.69 0.893 29.35 153.3
+Table VII presents the ablation study on the Degradation Embedding dimension ($D_{emb}$).
+Given that the Degradation Extractor serves as the sole source of degradation priors, the completeness and distinctiveness of the extracted information are paramount.
+To quantitatively evaluate the quality of the learned embeddings, we introduce two direct indicators: (1) Degradation Reconstruction PSNR (Rec.
+PSNR), calculated between the reconstructed degraded image and the target degraded input, which measures information integrity; and (2) Classification Accuracy (Cls.
+Acc.), which assesses the semantic separability of the degradation types.
+For the classification metric, we appended a simple MLP head to the frozen degradation extractor and trained it for only 1,000 iterations to map the embeddings to the seven degradation classes.
+As observed in Table VII, increasing $D_{emb}$ from 256 to 512 yields substantial gains: Rec.
+PSNR improves from 21.51 dB to 25.67 dB, and Cls.
+Acc.
+surges from 72.0% to 92.0%.
+This indicates that higher-dimensional embeddings effectively encode both the pixel-level details and the semantic categories of degradations.
+These improvements directly translate to the final restoration task, boosting AIO.
+PSNR to 27.69 dB.
+However, expanding to 1024 yields marginal returns despite higher costs.
+Thus, 512 is selected as the optimal balance.
+Table VIII presents the ablation study on global information supplementation.
 We compare our Prompt-C with Mamba’s bi-directional scanning scheme.
-Compared to Mamba’s bi-directional scan, our Prompt-C not only achieves slightly better performance but is also more efficient, reducing GFLOPs by nearly 37% (Table VI).
+Compared to Mamba’s bi-directional scan, our Prompt-C not only achieves slightly better performance but is also more efficient, reducing GFLOPs by nearly 37%.
 This highlights its effectiveness as a lightweight global information supplementation method.
-TABLE VI: Ablation Study on Different Methods for Global Information Supplementation.
-Method PSNR SSIM Params(M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt Matrix C (Ours) 27.69 0.893 29.35 153.3
+IV-F Generalization Capability and Failure Boundary Analysis
+To investigate the robustness of DPMambaIR beyond the training distribution, we conducted Out-of-Distribution (OOD) evaluations on two unseen degradation types: Raindrop and Pixelation.
+Importantly, these evaluations were performed by directly applying the pre-trained model to the test images without any additional training or fine-tuning.
+This rigorous setting serves to delineate the generalization boundaries of our Degradation-Aware Prompt mechanism.
+Robustness on Unseen Degradations.
+We evaluated the model on the Raindrop dataset[40] (58 images) and a synthetically generated Pixelated BSD68 dataset[29] (block size 2).
+As summarized in Table IX, DPMambaIR exhibits superior generalization, securing top-tier performance across all metrics.
+Specifically, on the Pixelation task, it achieves a PSNR of 23.35 dB, surpassing the second-best AdaIR by a significant margin of 1.96 dB.
+We attribute this robustness to our regression-based embedding strategy.
+Unlike static prompt approaches that risk model collapse on undefined inputs, DPMambaIR projects unseen degradations into a continuous latent space.
+This mechanism enables the model to approximate unknown corruption patterns within a learned manifold, preventing inference failure and yielding visually plausible results.
+A notable phenomenon is further observed in a specific low-light sample from the BSD68 dataset.
+When processing the pixelated version of this image, DPMambaIR automatically performs brightness enhancement alongside depixelation, capturing and correcting the underlying low-light degradation.
+In contrast, competing methods fail to address this compound issue.
+This capability underscores the potential of DPMambaIR in handling mixed degradations, suggesting that future work can further explore real-world restoration via more complex, fine-grained coupled degradation modeling.
+Failure Boundary Analysis.
+Despite these promising results, DPMambaIR is not without limitations.
+We observed a distinct performance bottleneck when processing images with heavy raindrops.
+As visually analyzed in Fig.
+10, while the model effectively removes light raindrops, it struggles with large, opaque water droplets that completely occlude the background.
+We attribute this to the physical nature of the degradation: removing heavy, opaque occlusions requires generative hallucination (inpainting) to synthesize missing content, whereas DPMambaIR operates on restoration cues.
+Future work could explore incorporating generative priors, such as Diffusion Models, to address such severe occlusions.
+TABLE IX: Quantitative comparison on unseen degradation types (OOD Testing). The best and second-best results are highlighted in bold and underline.
+Method Pixelation Raindrop
+(PSNR / SSIM / LPIPS) (PSNR / SSIM / LPIPS)
+MambaIR 20.80 / 0.7601 / 0.0568 19.29 / 0.7992 / 0.1051
+PromptIR 19.69 / 0.7479 / 0.0582 19.26 / 0.7775 / 0.1146
+IDR 20.64 / 0.7567 / 0.0494 17.40 / 0.7609 / 0.1276
+OneRestore 19.75 / 0.7415 / 0.0628 19.75 / 0.7792 / 0.1100
+AdaIR 21.39 / 0.7676 / 0.0512 19.07 / 0.7722 / 0.1225
+MoCEIR 19.89 / 0.7355 / 0.0544 21.46 / 0.8254 / 0.0919
+DPMambaIR 23.35 / 0.8003 / 0.0551 21.63 / 0.8337 / 0.0932
+Figure 10: Visual analysis of generalization and failure boundaries.
+(a) Success Cases: The top row demonstrates restoration on unseen Pixelation, where DPMambaIR concurrently removes artifacts and enhances inherent low-light details.
+The middle row shows effective removal of Light Raindrop.
+(b) Failure Case: The bottom row illustrates the boundary on Heavy Raindrop, where the model fails to hallucinate the background behind opaque occlusions.
 V Conclusion
@@
 This design enables the dynamic integration of degradation features into the state-space modeling process, allowing for adaptive handling of complex degradation scenarios while maintaining global degradation awareness.
-Furthermore, a lightweight High-frequency Enhancement Block (HEB) was introduced to complement the main framework by enhancing high-frequency details restoration with negligible computational overhead.
+Furthermore, a lightweight High-frequency Enhancement Block (HEB) is introduced to complement the main framework by enhancing high-frequency detail restoration with negligible computational overhead.
 This study underscores the importance of fine-grained degradation modeling and dynamic feature modulation in advancing All-in-One image restoration frameworks.